### v1

# Improved subsample-and-aggregate via the private modified winsorized mean

Abstract
We develop a univariate, differentially private mean estimator, called the private modified winsorized mean, designed to be used as the aggregator in subsample-and-aggregate. We demonstrate, via real data analysis, that common differentially private multivariate mean estimators may not perform well as the aggregator, even with a dataset with 8000 observations, motivating our developments. We show that the modified winsorized mean is minimax optimal for several, large classes of distributions, even under adversarial contamination. We also demonstrate that, empirically, the modified winsorized mean performs well compared to other private mean estimates. We consider the modified winsorized mean as the aggregator in subsample-and-aggregate, deriving a finite sample deviations bound for a subsample-and-aggregate estimate generated with the new aggregator. This result yields two important insights: (i) the optimal choice of subsamples depends on the bias of the estimator computed on the subsamples, and (ii) the rate of convergence of the subsample-and-aggregate estimator depends on the robustness of the estimator computed on the subsamples.
1 Introduction
Subsample-and-aggregate is a powerful algorithm that can produce a differentially privately version of any statistic [29]. For instance, through subsample-and-aggregate, complex machine learning models can be made private easily, with minimal additional programming. The algorithm is a two-step procedure, relying on a “subsample” step and an “aggregate” step. The idea, in brief, is to first split the data into disjoint subsets. Then, on each of these subsets, a nonprivate version of the desired statistic is computed. Next, the set of nonprivate statistics are combined, using a differentially private aggregator, producing a private estimate of the parameter of interest. Generally, if the mean of the subsamples is desired, the private aggregator is a basic differentially private clipped mean [29] or a widened winsorized mean [33]. However, to achieve pure differential privacy, both of these aggregators require tight input bounds on the statistic.
It is natural to consider the more recently developed, differentially private mean estimators, which have a weaker dependence, if any, on input bounds, as the private aggregator. However, currently, replacing the aggregator with any of the existing differentially private mean estimators comes with at least one of the following drawbacks: (i) only a weaker notion of differential privacy is satisfied, (ii) many subsamples are required for a stable mean estimate, resulting in large sample size requirements, (iii) the new algorithm has strong dependence on tuning parameters, (iv) strong assumptions on, or knowledge of, the population distribution is required, or (v) the estimator is not practically computable.
Here, we develop a new, univariate, differentially private mean estimator, called the private modified winsorized mean. Our estimator is designed to be used as the aggregator in subsample-and-aggregate and avoids these drawbacks. The private modified winsorized mean can satisfy either pure or zero-concentrated differential privacy, performs well in small samples or under adversarial contamination, has weak dependence on input bounds and other tuning parameters, and can be used to estimate the mean of a general distribution, where only two moments need to exist. The proposed estimator combines the winsorized mean of Lugosi and Mendelson, [28] and the private quantile estimation procedure of Durfee, [13].
1.1 Related work
Differentially private mean estimation has been well-studied. In the univariate setting, clipped and trimmed means have been the main approach [18, 24, 9, 23, 34, 13]. The clipped mean procedure operates by selecting an interval $(a,b)$, either empirically or a priori. Then, the dataset is projected onto the interval and an additive noise mechanism is used to ensure privacy. The trimmed mean functions similarly, removing extreme observations instead of projecting them. Existing estimators using these approaches have at least one of the previously mentioned drawbacks. In particular, Tsfadia et al., [34] considers approximately differentially private mean estimation. For stronger forms of privacy, many mean estimation algorithms require certain components of the population distribution as algorithm inputs, such as the number of moments possessed by the distribution [18, 23] or a lower bound on the variance of the population distribution [9]. On the other hand, algorithms aimed at being more “practical”, in the sense that they can be implemented without knowledge of the population distribution, lack minimax optimality or similar statistical guarantees [34, 13]. The most similar work is that of Bun and Steinke, [9], who develop a zero-concentrated differentially private trimmed mean estimate, combined with smooth sensitivity. Their estimator avoids many of the drawbacks listed, however, a lower bound on the variance is required, and it is not clear how to choose the smoothing parameter. In addition, their approach removes data from the estimator, which can be problematic at low sample sizes (see Section 4).
Multivariate estimators can be applied to univariate estimation problems. In the multivariate setting, approximate differentially private mean estimation has been well-studied [25, 5, 17, 12, 6, 11]. Under stronger privacy guarantees, many authors have considered mean estimators for a sample of subgaussian data [22, 7, 10, 4, 26]. Several authors have relaxed this assumption using moment conditions on the population distribution, through contamination models, or both [20, 35]. Unfortunately, some of these algorithms are not currently practically computable [20, 5, 26]. For the estimators that are computable, at small sample sizes ($n<1000$) we have observed empirically that they do not perform as well as univariate estimators applied coordinate-wise, and may be unstable or highly sensitive to tuning parameters (see Section 5). To use these estimators in subsample-and-aggregate, we require thousands of subsamples, and, consequentially, a very large sample size. This may not always be feasible.
Beyond mean estimation, several authors have considered improvements or variants of subsample-and-aggregate [33, 3, 21, 1]. Most of these focus on the “subsample” step of the algorithm [30, 3, 21]. Amin et al., [1] and Smith, [33] consider alternatives for the private aggregator. Smith, [33] show that using a private widened winsorized mean as the aggregator results in the same asymptotic distribution as the nonprivate estimator used in the subsample step. However, this aggregator is sensitive to input bounds, unless we relax the setting to approximate differential privacy. Amin et al., [1] consider an approximately differentially private, multivariate median aggregator in the setting of linear regression.
1.2 Organization and contributions
Our contributions are listed as follows: (i) We develop a new, univariate, differentially private mean estimator, designed to be used as the aggregator in subsample-and-aggregate. (ii) We show that this estimator is minimax optimal for several large classes of distributions, even under adversarial contamination (Theorem 3.1). As a corollary, we present a finite sample deviation bound for subsample-and-aggregate with the new aggregator (Corollary 5.1). (iii) We demonstrate that the estimator performs well relative to competitors in simulation (Section 4). (iv) We demonstrate via real data analysis that common differentially private multivariate mean estimators do not perform well as the aggregator, even with a dataset with 8000 observations (Section 5). (v) As a by-product of our analysis, we introduce a zero-concentrated differentially private version of the unbounded private quantile algorithm [13], as well as some concentration inequalities for these quantile estimates (Lemma D.1 and Lemma E.1E.2). In addition, we correct an error in the proof of Theorem 1 of Lugosi and Mendelson, [28], see Appendix F.
The paper is organized as follows. Section 2 presents the necessary background on differential privacy and mean estimation. Section 3 presents the private modified winsorized mean, our main theoretical result, and a more practical version of the private modified winsorized mean. Section 4 presents a simulation study, comparing the proposed estimator with various existing estimators. Section 5 presents a theoretical and empirical analysis of the improved subsample-and-aggregate algorithm.
2 Preliminary material
2.1 Problem statement
First, we introduce notation to formalize the problem. For a positive integer $m$, let $[m]=\{1,\ldots,m\}$. Suppose that we have access to a random sample of size $2n$, of which at most $2\eta n$ points, for $0\leq\eta\leq 1$, are corrupted in that they have been changed to arbitrary values. The corrupted sample is denoted $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$. The uncorrupted points, denoted $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, are assumed to be an independent, identically distributed sample from a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$. The aim is to estimate $\mu$ while maintaining differential privacy (to be defined in Section 2.2). This setup is summarized as Condition 1.
Condition 1.
The uncorrupted points, $X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}$, are an independent, identically distributed sample from a univariate distribution $F$ with mean $\mu\in\mathbb{R}$ and variance $\sigma^{2}<\infty$. The points $X_{1}^{\prime},\ldots,X_{n}^{\prime},Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$ are equal to the uncorrupted points at all but $2\eta n$ points, $0\leq\eta\leq 1$, which have been set to arbitrary, unknown values.
2.2 Differential privacy
Next, we introduce the essential concepts from differential privacy used in this work [15, 16]. Define a dataset of size $n\in\mathbb{N}$ to be a set of $n$ real numbers, and let $\mathcal{D}_{n}$ be the set of datasets of size $n$. We say that a dataset $\mathbf{x}_{n}\in\mathcal{D}_{n}$, is adjacent to another dataset $\mathbf{y}_{n}\in\mathcal{D}_{n}$ if $\mathbf{x}_{n}$ and $\mathbf{y}_{n}$ differ by exactly one point. Let $\mathcal{A}_{n}$ be the set of pairs of adjacent datasets of size $n$. Next, let $G_{\mathbf{x}_{n}}$ be a probability distribution over $\mathbb{R}^{d}$ with $d\in\mathbb{N}$ that depends on $\mathbf{x}_{n}$. That is, $G_{.}\colon\mathcal{D}_{n}\to\mathcal{M}_{1}(\mathbb{R}^{d}),$ where $\mathcal{M}_{1}(\mathbb{R}^{d})$ is the set of (Borel) probability distributions on $\mathbb{R}^{d}$. Assuming that $G_{\mathbf{x}_{n}}$ is absolutely continuous, let $g_{\mathbf{x}_{n}}$ be the associated density. We can now define (pure) differential privacy (PDP).
Definition 2.1.
The quantity $\theta\sim G_{\mathbf{x}_{n}}$ is $\varepsilon$-differentially private ($\varepsilon$-PDP) if
$\sup_{(\mathbf{y}_{n},\mathbf{z}_{n})\in\mathcal{A}_{n}}\sup_{x\in\mathbb{R}}\frac{g_{\mathbf{y}_{n}}(x)}{g_{\mathbf{z}_{n}}(x)}\leq e^{\varepsilon}.$
Here, $\theta$ is a $d$-dimensional differentially private quantity and $\varepsilon$ is the privacy budget, where values of $\varepsilon$ closer to zero imply higher levels of privacy. In some settings, pure differential privacy can be too strict. A relaxed version of differential privacy is zero-concentrated differential privacy (zCDP) [8]. For $\alpha>1$, let ${\rm D}_{\alpha}(F|G)$ denote the $\alpha$-Renyi Divergence between two distributions, $F$ and $G$.
Definition 2.2.
Given $\rho>0$, the quantity $\theta\sim F_{\mathbf{x}_{n}}$ is $\rho$-zero-concentrated differentially private ($\rho$-zCDP) if, for all $\alpha\in(1,\infty)$,
$\sup_{(\mathbf{y}_{n},\mathbf{z}_{n})\in\mathcal{A}_{n}}{\rm D}_{\alpha}(G_{\mathbf{y}_{n}}|G_{\mathbf{z}_{n}})\leq\alpha\rho.$
Similar to $\epsilon$ in PDP, $\rho$ is the privacy budget for zCDP, where values of $\rho$ closer to zero imply higher levels of privacy. Quantities that are $\varepsilon$-PDP are also $\varepsilon^{2}/2$-zCDP quantities. The following two properties are useful for developing differentially private estimates.
Proposition 2.1.
The following hold:
  • Post-processing: For any function $f$ that does not depend on $\mathbf{x}_{n}$, defined on a subset of $\mathbb{R}^{d}$, if $\theta$ satisfies $\varepsilon$-PDP ($\rho$-zCDP), then $f(\theta)$ satisfies $\varepsilon$-PDP ($\rho$-zCDP).
  • Composition: If $\theta_{1},\theta_{2}$ satisfy $\varepsilon_{1}$-PDP and $\varepsilon_{2}$-PDP ($\rho_{1}$-zCDP and $\rho_{2}$-zCDP), respectively, then $(\theta_{1},\theta_{2})$ satisfies $(\varepsilon_{1}+\varepsilon_{2})$-PDP ($(\rho_{1}+\rho_{2})$-zCDP).
Proposition 2.1 first states that functions of differentially private quantities are still private. Additionally, it states that releasing two differentially private quantities results in a jointly differentially private estimate, with a larger privacy parameter.
Next, we introduce the simplest way to generate a differentially private statistic: additive noise mechanisms. For a univariate statistic $T\colon\mathcal{D}_{n}\to\mathbb{R}$, define the global sensitivity of $T$ as
$\mathrm{GS}(T)=\sup_{(\mathbf{y}_{n},\mathbf{z}_{n})\in\mathcal{A}_{n}}|T(\mathbf{y}_{n})-T(\mathbf{z}_{n})|.$
We have the following proposition, introducing the Laplace and Gaussian mechanisms.
Proposition 2.2.
Let $T\colon\mathcal{D}_{n}\to\mathbb{R}$ be any statistic with $\mathrm{GS}(T)<\infty$. If $Z$ is a standard Laplace random variable, then
$T(\mathbf{x}_{n})+\mathrm{GS}(T)Z/\varepsilon$
is $\varepsilon$-differentially private. This is referred to as the Laplace mechanism. If $Z$ is a standard Gaussian random variable, then
$T(\mathbf{x}_{n})+\mathrm{GS}(T)Z/\sqrt{2\rho}$
is $\rho$-zCDP. This is referred to as the Gaussian mechanism.
Proposition 2.2 illustrates how correctly calibrated noise can be added to a statistic to ensure differential privacy.
2.3 Lugosi and Mendelson’s modified winsorized mean
We now review the non-private inspiration for the proposed estimator. Lugosi and Mendelson, [28] recently showed that, out of the class of non-private estimators, a modified winsorized mean estimator is minimax optimal, for the model given in Condition 1. Given a trimming proportion $20<p<1/$ 2, the idea is to first use half of the sample, $\mathbf{Y}_{n}^{\prime}=\{Y_{i}^{\prime}\}_{i=1}^{n}$, to estimate the $p$th quantile and the $(1-p)$th quantile of $F$. We denote these quantile estimates by $\hat{\xi}_{p}$ and $\hat{\xi}_{1-p}$, respectively. Then, we project the other half of the observations $\mathbf{X}_{n}^{\prime}=\{X_{1}^{\prime},\ldots,X_{n}^{\prime}\}$ onto the interval $[\hat{\xi}_{p},\hat{\xi}_{1-p}]$. The sample mean of the projected observations is taken as the final estimate of $\mu$. Formally, for $0<q\leq 1$, let $\hat{\xi}_{q}$ be the (left-continuous) $q$th quantile of the empirical distribution defined by $\mathbf{Y}_{n}^{\prime}$. For $a\leq b$, define
$\phi_{a,b}(x)=\begin{cases}a&x<a\\ x&a\leq x\leq b\\ b&x>b\end{cases}.$
The non-private trimmed mean is
$\hat{\mu}_{p}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\hat{\xi}_{p},\hat{\xi}_{1-p}}(X_{i}^{\prime}).$
Lugosi and Mendelson, [28] prove an upper bound on $|\hat{\mu}_{p}-\mu|$ that holds with probability $1-\delta$, with $\delta$ at most exponentially small in $n$. Their result shows that taking $p=8\eta+12\log(4/\delta)/n$ yields an estimator that is minimax optimal. Here, we extend $\hat{\mu}_{p}$ to the differentially private setting. The motivation is that, for a carefully chosen $p$, the optimality in the nonprivate setting should translate to the private setting, provided we are able to estimate extreme quantiles well under the restriction of differential privacy.
2.4 Private quantile estimation
A private version of the trimmed mean, $\hat{\mu}_{p}$, requires differentially private estimation of extreme quantiles. We make use of the unbounded private quantile estimator, developed by Durfee, [13] and studied briefly by Ramsay and Diaz-Rodriguez, [31]. We chose the unbounded quantile estimator as it has been shown to estimate extreme quantiles consistently, whereas other private quantile estimators do not [31]. Specifically, when there is no contamination, our mean estimator requires a private analogue of an $O(1/n)$ quantile. It is expected that a private $O(1/n)$th quantile would be consistent for the infimum of the support of $F$, as $n\to\infty$. [31] show that other private quantile estimators do not have this property, however, the unbounded estimator does.
We present the modified estimator given by Ramsay and Diaz-Rodriguez, [31]. Let $V,V_{1},V_{2},\ldots$ be a sequence of independent, standard exponential random variables. Suppose that we wish to estimate the $q$th quantile of the population distribution privately, where the desired quantile is an upper quantile, (i.e., $1/2\leq q\leq 1$). First, we supply the algorithm with the following input parameters: a grid size parameter $\beta>1$, a lower bound on the desired quantile $\ell$, a dataset with empirical distribution $F_{n}$, and two privacy budgets $\varepsilon_{1},\varepsilon_{2}>0$. Given these parameters, the algorithm proceeds as follows:
  1. 1.
    Compute $\hat{q}=q+V/n\varepsilon_{1}$ and set $i=1$.
  2. 2.
    Increase $i$ by one until $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\varepsilon_{2}>\hat{q}$. That is, stop at the first $i$ such that $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\varepsilon_{2}>\hat{q}$.
  3. 3.
    Output $\beta^{i}+\ell-1$.
If $20<q<1/$ 2, we supply an upper bound on the desired quantile, $u$. Then, we negate the dataset, and run the upper-quantile algorithm, with input quantile $1-q$ and $\ell=-u$. This produces a negated estimate of the lower quantile that we negate to get the final quantile estimate. We refer to these quantile estimates by $\tilde{\xi}_{q}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}(\varepsilon_{1},\varepsilon_{2},\ell,u,\beta)$. The algorithm is $(\varepsilon_{1}+\varepsilon_{2})$-differentially private [13].
We show (Lemma D.1) that replacing $\varepsilon_{i}$ with $\sqrt{\rho_{i}}$ and taking $V,V_{1},V_{2},\ldots$ to be standard Gaussian random variables produces $(\rho_{1}+\rho_{2})$-zero-concentrated differentially private quantiles. We denote the estimates generated by the zero-concentrated differentially private version of the algorithm by $\tilde{\xi}_{q}^{\prime}\coloneqq\tilde{\xi}_{q,\mathbf{X}_{n}}^{\prime}(\rho_{1},\rho_{2},\ell,u,\beta)$. As a by-product of our analysis, we prove a simple concentration result concerning both versions of these private quantiles (Lemmas E.1-E.2).
3 A new differentially private mean estimator
3.1 Definition
We now introduce our private mean estimator, the private modified winsorized mean, drawing inspiration from Lugosi and Mendelson, [28] and Durfee, [13]. Given a trimming proportion $20<p<1/$ 2 and private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2}>0$ (PDP) or $\rho_{1},\rho_{2}>0$ (zCDP), we first compute $\tilde{\xi}_{p}$ and $\tilde{\xi}_{1-p}$ (PDP) or $\tilde{\xi}_{p}^{\prime}$ and $\tilde{\xi}_{1-p}^{\prime}$ (zCDP). Recall that these are the private estimates of the extreme quantiles, based on $\mathbf{Y}_{n}^{\prime}$. The data $\mathbf{X}_{n}^{\prime}$ are then projected onto the interval $[\tilde{\xi}_{p},\tilde{\xi}_{1-p}]$ (PDP) or $[\tilde{\xi}_{p}^{\prime},\tilde{\xi}_{1-p}^{\prime}]$ (zCDP). Lastly, we take the sample mean of the projected observations, and add calibrated noise via the Laplace mechanism (PDP) or the Gaussian mechanism (zCDP). Given $14e^{-n}<\delta<$ 1, our main result is a bound on the deviations of the private modified winsorized mean about $\mu$ that holds with probability $1-\delta$. Under two mild conditions, taking
$p=\zeta\coloneqq\zeta(n,\eta,\delta,\ell,u,\beta)=16\eta+\frac{112}{3}\frac{\log(32(\beta(u-\ell)/(\beta-1)\vee 1)/\delta)}{n},$
for both PDP and zCDP, yields minimax optimal estimates of $\mu$ for the problem stated in Condition 1, where the class of estimators is constrained to those that are differentially private (PDP or zCDP). Here, for two real numbers $a$ and $b$, we write $a\wedge b$ ($a\vee b$) to denote the minimum (maximum) of $a$ and $b$. We formally define the estimators as follows.
Definition 3.1.
Given private quantile parameters $\ell,u\in\mathbb{R}$, $\beta>1$, and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}>0$ (PDP) or $\rho_{1},\rho_{2},\rho_{3}>0$ (zCDP), the private modified winsorized means are
$\tilde{\mu}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\tilde{\xi}_{\zeta},\tilde{\xi}_{1-\zeta}}(X_{i}^{\prime})+Z_{1}\frac{(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}}\qquad\text{and}\qquad\tilde{\mu}^{\prime}=\sum_{i=1}^{n}\frac{1}{n}\phi_{\tilde{\xi}_{\zeta}^{\prime},\tilde{\xi}_{1-\zeta}^{\prime}}(X_{i}^{\prime})+Z_{2}\frac{(\tilde{\xi}_{1-\zeta}^{\prime}-\tilde{\xi}_{\zeta}^{\prime})}{n\sqrt{2\rho_{3}}},$
where $Z_{1}$ is a standard Laplace random variable and $Z_{2}$ is a standard Gaussian random variable.
It follows from basic composition (Proposition 2.1) that $\tilde{\mu}$ is $(2\varepsilon_{1}+2\varepsilon_{2}+\varepsilon_{3})$-differentially private and that $\tilde{\mu}^{\prime}$ is $(2\rho_{1}+2\rho_{2}+\rho_{3})$-differentially private. Our main result requires two conditions on $\zeta$, $F$, and the input parameters of the private quantile algorithm $(\ell,u,\beta)$.
Condition 2.
It holds that $\ell\leq\xi_{5\zeta/4}$, $u\geq\xi_{1-5\zeta/4}$, and $\zeta<1/2$.
Condition 2 requires that the interval $(\ell,u)$ contains the interval $(\xi_{q/4},\xi_{1-q/4})$, which, colloquially, means “most of the uncontaminated data”. Our theoretical results show that the bounds $\ell\leq\xi_{q/4}$, $u\geq\xi_{1-q/4}$ can be loose, replaced by loose bounds on $\mu$ and $\sigma$, or replaced with a weaker condition, see Remark 3.3 and Remark 3.4.
Condition 3.
It holds that
$1<\beta<1+\left[\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}\wedge\frac{\xi_{1-3\zeta/4}-\xi_{1-5\zeta/4}}{\xi_{1-5\zeta/4}-\ell+1}\right]\coloneqq 1+b_{n}.$
Condition 3 says that the grid size used for quantile estimation cannot be too coarse. Combining this condition with our main result yields a range of $\beta$ values that produce minimax optimal estimates, up to logarithmic terms, see Remark 3.2. We now introduce our main theoretical result. For quantities $a,b$, we write $a\lesssim b$ ($a\gtrsim b$) when there is a universal constant $C>0$ such that $a\leq Cb$ ($a\geq Cb$).
Theorem 3.1.
Take $\eta\geq 0$, $14e^{-n}<\delta<$ 1, $\varepsilon_{1},\varepsilon_{2}>112/3$, $\rho_{1},\rho_{2}>112/3$ and $\varepsilon_{3},\rho_{3}>0$. For all $n\geq 1$, $u,\ell\in\mathbb{R}$, and $\beta>1$ such that Conditions 13 hold, with probability at least $1-\delta,$
$|\tilde{\mu}-\mu|\lesssim\sigma\left(\sqrt{\eta}\vee\sqrt{\frac{\log(u-\ell)+\log(\beta/(\beta-1))+\log(4/\delta)}{n}}\vee\frac{1}{\sqrt{n}\varepsilon_{3}}\right),$ (3.1)
and
$|\tilde{\mu}^{\prime}-\mu|\lesssim\sigma\left(\sqrt{\eta}\vee\sqrt{\frac{\log(u-\ell)+\log(\beta/(\beta-1))+\log(4/\delta)}{n}}\vee\frac{1}{\sqrt{n\rho_{3}}}\right).$ (3.2)
The proof of Theorem 3.1 is in Appendix C. To prove Theorem 3.1 we first prove a concentration inequality on the private quantiles. As a result, it suffices to bound the quantity $n^{-1}\sum_{i=1}^{n}\phi_{\xi_{\zeta},\xi_{1-\zeta}}(X_{i}^{\prime})-\mu$. For this task, the proof is similar to the proof of Theorem 1 of Lugosi and Mendelson, [28]. However, their proof contains an error, that is resolved here, see Appendix F.
Interpreting Theorem 3.1, we first observe that, provided that $\log(\beta(u-\ell)/(\beta-1))\lesssim\log(4/\delta)$, and Conditions 2 and 3 are satisfied, the rate given in Theorem 3.1 is minimax optimal, regardless of the level of contamination. Observe that the upper bounds in Theorem 3.1 each have three terms: the contamination cost, the sampling error cost, and the privacy cost, listed in order of appearance. The contamination cost and the sampling error cost, together, match the minimax lower bound in the non-private setting [28]. Furthermore, for the PDP mean, the sampling error cost and the privacy cost match the minimax lower bound in the private setting [23]. Using these results together, the private modified winsorized means are minimax optimal.
To see this, note that when there is no contamination, $\eta=0$, the rates match the established minimax bounds for PDP mean estimation [23]. Neither of the components that remain with $\eta=0$ depend on $\eta$, and as such, their contribution to the rate does not change based on the level of contamination. Moreover, ignoring the privacy error, the rates match the established minimax optimal rates for non-private estimation, at any level of contamination [28]. As a result, the PDP modified winsorized means then has to be minimax optimal for all levels of contamination, provided $\eta$ is such that Conditions 2 and 3 are satisfied. Finally, given that PDP lower bounds imply zCDP lower bounds, we have that the zCDP modified winsorized mean is also minimax optimal for this problem.
Remark 3.2 (On choosing $\beta$).
For minimax optimality, we require $\log(\beta/(\beta-1))\lesssim\log(4/\delta)$. Combining this with Condition 3 yields a recommended range for the grid size: $(1-(\delta/4)^{s})^{-1}\leq 1+(\delta/4)^{s}/(1-(\delta/4)^{s})\leq\beta\leq 1+b_{n}$, for some $s>0$. For example, we can select $\beta=O(1+n^{-s}/(1-n^{-s}))$, provided $\delta<n^{-s}<b_{n}$ for some $s>1$. This ensures that the bound (3.1) holds with probability at least $1-n^{-s}$. Regarding $b_{n}$, bounded support makes for a smaller $b_{n}$, while unbounded support, with heavy tails, makes for a larger $b_{n}$. From the proof of Theorem 3.1, we only need the private quantiles to be in some range of extreme quantiles of the population distribution $(\xi_{\zeta/4},\xi_{4\zeta})$ for the lower quantile and $(\xi_{1-4\zeta},\xi_{1-\zeta/4})$ for the upper quantile. This range is large for distributions with unbounded support, and small for those with bounded support. For instance, for the standard uniform distribution, $b_{n}=O(1/n)$, for the standard exponential distribution, $b_{n}\gtrsim 1/(u-\log n+1)$, and for Student’s $t$-distribution with three degrees of freedom, $b_{n}\gtrsim 1,$ when $u=2\xi_{1-5\zeta/4}$ and $\ell=-u$. Thus, we must take a finer grid when applying the procedure to distributions with bounded support, but a coarse grid is acceptable for distributions with unbounded support.
Remark 3.3 (On choosing $u$ and $\ell$).
Condition 2 requires that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. Then, (3.1) and (3.2) imply that taking $\log(u-\ell)\lesssim\log(1/\delta)$ produces a modified winsorized mean that is minimax optimal. Choosing $\ell,u$ thus amounts to ensuring these two bounds are satisfied. First, observe that, in general, we take $\delta$ to be small, say, $\delta=n^{-k}$ for some $k>0$. The bound $\log(u-\ell)\lesssim\log(1/\delta)$ then requires that $u-\ell$ is at most polynomial in $n$. Now, we need $u,\ell$ such that $\ell<\xi_{5\zeta/4}<\xi_{1-5\zeta/4}<u$. For $F$ with bounded support, this is straightforward. If the support is unbounded, we can specify $u,\ell$ according to the following argument. We have that, by Chebyshev’s inequality, $5\zeta/4\leq\sigma^{2}/(\xi_{1-5\zeta/4}-\mu)^{2}$. This implies that $\xi_{1-5\zeta/4}\lesssim\sigma/\zeta^{1/2}+\mu$. Therefore, Condition 2 is implied by $u\gtrsim\sigma/\zeta^{1/2}\vee\mu$. Then, using $\delta=n^{-k}$, $\sigma/\zeta^{1/2}\lesssim\sigma\sqrt{n/k\log n}.$ Assuming that $\mu,\sigma$ are constant in $n$, we take $u\geq n^{1/2+r},$ $\ell\leq-n^{1/2+r},$ for some $r>0$. For large $n$, this implies Condition 2. If one has instead (potentially very loose) bounds such that $-\mu_{0}<\mu<\mu_{0}$ and $\sigma<\sigma_{0}$, then taking $u\gtrsim\sqrt{n\sigma_{0}}\vee\mu_{0}$ and $\ell\lesssim-(\sqrt{n\sigma_{0}}\vee\mu_{0})$ implies Condition 2, and ensures that $u-\ell$ is at most polynomial in $n$.
Remark 3.4 (Weakening Condition 2).
Condition 2 requires that the interval $(\ell,u)$ contains most of the uncontaminated data. This condition can be relaxed to only require that $u>\xi_{5\zeta/4}$ and $\ell<\xi_{1-5\zeta/4}$, (i.e., $u$ is an upper bound on the lower quantile and $\ell$ is a lower bound on the upper quantile), which is weaker. This will come at the expense of replacing $u-\ell$ in (3.1) and (3.2) with $\log n$. This modification still constitutes a minimax optimal rate, up to logarithmic terms. However, given that the bounds on the data can be very loose, as per (3.1) and (3.2), and the results in Section 4, we leave Condition 2 unchanged.
Remark 3.5 (Subgaussian distributions and more).
As usual, (e.g., as in the work of Lugosi and Mendelson, [28]), we can derive stronger bounds under stronger assumptions. Concretely, we use Chebyshev’s inequality to generate the first and last terms in (3.1) and (3.2). This can be replaced with the tightest tail bound given the assumptions. For instance, if we instead assume the uncontaminated data are subgaussian observations, we use a subgaussian tail bound to get tighter deviation bounds. In that case, the private trimmed mean also achieves optimal rates.
Remark 3.6 (Multivariate case).
The private trimmed mean can be extended in a coordinate-wise fashion to give a simple definition of a multivariate estimator. Given the coordinate-wise nature, this estimator will not achieve minimax optimality. The resulting bound contains a term $\sqrt{d\eta}$. There are (non-private) estimators (e.g., in Lugosi and Mendelson, [28]) which depend on the contamination only through the term $\sqrt{\eta\lambda_{1}}$, where $\lambda_{1}$ is the largest eigenvalue of the covariance matrix. It may be possible to extend a similar logic as in the univariate case to this more efficient multivariate estimator. However, even non-privately, this is a difficult problem computationally [27], and the goal of this work was to develop a practical and fast algorithm to be used as part of subsample-and-aggregate, not to develop a new, minimax optimal multivariate estimator. As such, we leave extensions of the proposed estimator to the multivariate setting for future work.
We close this section by stating the sample complexity analogue of Theorem 3.1, when $\eta=0$, which may be of interest to some readers. Note, a version of Corollary 3.7 at positive corruption values $\eta$ that are constant in $n$ is impossible, since the minimax lower bound stipulates that no estimator can be consistent, at least without imposing additional restrictions on $F$.
Corollary 3.7.
Given $14e^{-n}<\delta<$ 1, and $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\rho_{1},\rho_{2},\rho_{3}>0$, if the conditions of Theorem 3.1 hold and $\eta=0$, then there exists a universal constant $K>0$ such that for all $t>0$, $|\tilde{\mu}-\mu|\leq t$ with probability at least $1-\delta,$ provided
$n\geq K\left(\sigma^{2}\frac{\log(u-\ell)-\log((\beta-1)/\beta)+\log(4/\delta)}{t^{2}}\vee\frac{1}{t^{2}\varepsilon_{3}}\right)$ (3.3)
and $|\tilde{\mu}^{\prime}-\mu|\leq t$ with probability at least $1-\delta,$ provided
$n\geq K\left(\sigma^{2}\frac{\log(u-\ell)-\log((\beta-1)/\beta)+\log(4/\delta)}{t^{2}}\vee\frac{1}{t^{2}\sqrt{\rho_{3}}}\right).$ (3.4)
3.2 A more practical private clipped mean
Next, we present adjustments to the proposed estimator $\tilde{\mu}^{\prime}$ that we recommend in practice. First, note that splitting the sample in half to compute each of the two steps of the estimator does not have any algorithmic or intuitive justification. It is an arbitrary choice that simplifies the proof of the deviation bound. Therefore, we modify the algorithm and use all data in each estimation step. Note that this does not change the privacy guarantees of the estimators.
Next, observe that no effort was made to optimize the constants in the clipping parameter $\zeta$ in the theoretical exposition of Section 3. Therefore, using $\zeta$ as a guide, we present the following practical version of the clipping parameter. The user is asked to select two parameters, $0<C<0.025n$ and $20\leq\eta<1/$ 2. As before, $\eta$ is the maximum expected level of contamination. The new parameter, $C$, represents the number of observations the user would like to trim from either end, if the contamination level was 0. We cap $C$ at $0.025n$ as otherwise, at very small sample sizes, $C$ may be larger than $n/2$. Finally, the clipping parameter is computed to be $C/n\vee\eta$, and is used in place of $\zeta$. We use a simulation study to investigate different values of $C$ and $\eta$, and found that the value of $C$ was not particularly important (see Appendix A.3). In addition, as expected, smaller $\eta$ was better for uncontaminated, light-tailed, uncontaminated data (Gaussian, Exponential and Gaussian mixtures) but larger $\eta$ was preferable for heavy tailed or contaminated data (Contaminated Gaussian and student $t$) (see Appendix A.3).
Next, not covered by the present theory, is how to best split the total budget $\rho$ amongst $\rho_{1},\rho_{2},$ and $\rho_{3}$. For this, we simulated the performance of the estimator under three scenarios: more weight to the quantile estimation, equal weight to the mean estimation and quantile estimation, and more weight to the mean estimation. We found that more weight to the mean estimation, specifically setting $\rho_{1}=\rho_{2}=\rho/16$ and, $\rho_{3}=3\rho/4$ worked best. Thus, we split the budget as such in the following results (see Appendix A.4). We abuse the notation and call this practical estimator $\tilde{\mu}^{\prime}$.
4 Simulation study
Next, we compare the empirical performance of the practical estimator, denoted $\tilde{\mu}^{\prime}$, as described in the preceding subsection, to several existing estimators. In this section, for brevity, we only consider zCDP estimators. For each run of the simulation for $\tilde{\mu}^{\prime}$, we set $\eta=0$ or $\eta=0.3$, take $C$ to be uniformly, randomly chosen between 1 and 100, and set $\beta=1.001$ (as in [13]), $\rho_{1}=\rho_{2}=\rho/16$, $\rho_{3}=3\rho/4$, and $\ell=-50<u=50$. For the existing estimators, any estimator that required input bounds had input bounds set at $\ell=-50<u=50$, and for any estimator that required setting number of moments, we set the number of moments at $2$. We compared our algorithm to the estimators of Foygel Barber and Duchi, [18], denoted Duchi with $r=25$, Kamath et al., [22] denoted Kamath with $\alpha=1$, $R=u=50$, and $\beta=0.1$, and the Laplace Log-Normal mean algorithm of Bun and Steinke, [9] denoted Bun with $m=C$ and either $t=0.01$ or $t=1$. In addition, for Kamath, we set $m=5\log 20$ instead of $m=200\log 20$ to accommodate the small sample sizes. We consider the performance of the estimators under five different population distributions (Gaussian, Gaussian mixture, Exponential, heavy tailed Student’s $t$, and a 20%-contaminated Gaussian) under varying privacy budgets $\rho\in\{1,10,100\}$ and sample sizes $n\in\{50,100,500,1000\}$. Each combination of the above scenarios was run 250 times.
Figure 1: The empirical $\log MSE$ of the proposed, practical mean estimator, $\tilde{\mu}$, with $\eta=0$ and $\eta=1$, compared to the estimator of Bun and Steinke, [9], Bun, with $t=1$, compared across a variety of different population distributions (Gaussian, mixture of Gaussians, Skewed, Heavy-Tailed, and a Contaminated Gaussian) and different privacy budgets ($\rho$).
Several of the results are omitted from the main figure to save space. The complete tables of all results are available in the Appendix A.2. For the main results, we note that Duchi and Kamath did not perform as well as Bun and $\tilde{\mu}$, so we omit the results from those estimators. Additionally, for Bun, $t=1$ performed better, so we omit the results with $t=0.01$.
Figure 1 displays the log-transformed empirical mean squared errors of each estimator, for each simulation scenario. At high levels of privacy and small sample sizes, $\tilde{\mu}^{\prime}$ outperforms Bun at both levels of $\eta$. To explain this, we explore the empirical mean squared error of the estimated means of Bun and $\tilde{\mu}$ before the noise is added, and the empirical variance of the additive noise for each estimator, recorded in Tables 1 and 2, respectively. These convey the difference between using a clipped mean and a trimmed mean, and the difference in the scale and type of noise added. Table 1 shows that the MSE related to the clipped mean is similar to that of the trimmed mean for small sample sizes, and smaller for large sample sizes. On the other hand, Table 2 shows that the average magnitude of noise added to construct $\tilde{\mu}$ is almost always smaller than that used to construct Bun at high levels of privacy. This explains why, at high levels of privacy and small sample sizes, $\tilde{\mu}^{\prime}$ outperforms Bun for both levels of $\eta$.
For the light-tailed, uncontaminated data (Gaussian, Exponential and Gaussian mixtures) $\tilde{\mu}^{\prime}$ with $\eta=0$ performs the best out of the three estimators. Intuitively, trimming less is better for the light-tailed, uncontaminated data. This can explain the fact that $\tilde{\mu}^{\prime}$ with $\eta=0$ outperforms $\tilde{\mu}^{\prime}$ with $\eta=0.3$. The differences between Bun and $\tilde{\mu}^{\prime}$ with $\eta=0$ can again be explained with Tables 1 and 2. The empirical mean squared error of the clipped mean is almost always lower than that of the trimmed mean. At high levels of privacy and large sample sizes, the variance of noise added to the non-private estimates is lower for that of Bun, but this effect is small compared to the differences between the (non-private) clipped mean and the (non-private) trimmed mean.
Table 1: Empirical mean squared errors, over 250 runs, of the trimmed mean, Bun, with $t=1$, and the clipped mean, $\tilde{\mu}$, with $\eta=0$, before the noise is added, compared across various population distributions. Here, $C$ and $m$ (as defined in [9]) were randomly selected between 1 and 100.
Gaussian Gaussian Mixture Skewed
$n$ Bun $\tilde{\mu}^{\prime},\ \eta=0$ Bun $\tilde{\mu}^{\prime},\ \eta=0$ Bun $\tilde{\mu}^{\prime},\ \eta=0$
50 0.0201 0.0222 0.5333 0.5517 0.0250 0.0297
100 0.0105 0.0107 0.3132 0.2889 0.0147 0.0130
500 0.0025 0.0025 0.0507 0.0464 0.0072 0.0024
1000 0.0013 0.0013 0.0284 0.0258 0.0068 0.0014
For skewed distributions, the differences in performance are greater. When data are skewed, trimming a constant fraction of data when there is no contamination, (i.e., $\eta=0.3$) results in higher error estimates. This is because clipping a constant fraction of estimates will result in an inconsistent estimator, non-privately. A similar phenomenon occurs with the Bun estimator. To elaborate, observations in the algorithm of Bun are removed, rather than projected. Removing an equal number of observations from each extreme end of the dataset is predicated upon an implicit symmetry assumption. When the distribution is skewed, and we remove $m$ observations from either end of the sample, we are trimming a tail on one side and not on the other, resulting in a biased estimator. This can also be seen theoretically, where the (non-private) trimmed mean is inconsistent whenever $m/n\to p>0$ (see Lemma G.1).
On the other hand, when the trimmed observations are contaminants, it is more advantageous to remove them, rather than use their projected versions. That is why we see, in the contaminated Gaussian, Bun often outperforms $\tilde{\mu}^{\prime}$ with $\eta=0$. Unsurprisingly, $\tilde{\mu}^{\prime}$ with $\eta=0.3$ performs best under heavy tails and contamination. This highlights the robustness of the estimator when $\eta>0$. In general, we recommend a moderate value of $\eta$ if no prior knowledge of the dataset is given.
Table 2: Empirical variance of the additive noise in the algorithms, over 250 runs, of the trimmed mean, Bun, with $t=1$, and the clipped mean, $\tilde{\mu}$, with $\eta=0$, compared across various population distributions. Here, $C$ and $m$ (as defined in [9]) were randomly selected between 1 and 100.
Gaussian Gaussian Mixture Skewed
$\rho$ $n$ Bun $\tilde{\mu}^{\prime},\ \eta=0$ Bun $\tilde{\mu}^{\prime},\ \eta=0$ Bun $\tilde{\mu}^{\prime},\ \eta=0$
1 50 1.46e+01 3.81e-03 1.50e+01 3.68e-02 1.51e+01 4.69e-03
100 9.47e-01 8.71e-04 1.32e+00 1.13e-02 9.74e-01 7.32e-04
500 1.46e-02 4.83e-05 2.50e-02 4.89e-04 1.51e-02 3.88e-05
1000 3.39e-03 1.38e-05 8.05e-03 1.38e-04 3.18e-03 9.37e-06
100 50 4.72e-04 4.44e-05 4.85e-04 5.33e-04 4.79e-04 3.62e-05
100 2.66e-05 1.10e-05 3.66e-05 1.17e-04 2.70e-05 9.76e-06
500 2.83e-07 5.17e-07 7.39e-07 4.78e-06 2.84e-07 5.09e-07
1000 5.14e-08 1.30e-07 1.89e-07 1.20e-06 4.92e-08 1.20e-07
5 Improved subsample-and-aggregate
We now demonstrate that the proposed mean estimator can be used to improve subsample-and-aggregate. Subsample-and-aggregate, developed by Nissim et al., [29], and extended by Smith, [33], is a method for estimating virtually any quantity privately. Define a multivariate dataset to be a set of $n$, $p$-dimensional vectors, and let $\mathcal{D}_{n,p}$ be the set of all such datasets. Consider a multivariate dataset, say $\mathbf{X}_{n}$, a set of $n$ realizations of $p$-dimensional random vectors, that are independent and identically distributed. To compute a statistic $T\colon\mathcal{D}_{n,p}\to\mathbb{R}^{d}$ for $d\geq 1$, (e.g., a summary statistic or a set of parameters for a machine learning model), from $\mathbf{X}_{n}$, privately, we can:
  1. 1.
    Split the dataset $\mathbf{X}_{n}$ into $m$ disjoint groups of size $k=\lfloor m/n\rfloor$, say $\mathbf{X}_{n}^{(1)},\ldots,\mathbf{X}_{n}^{(m)}$.
  2. 2.
    Compute $T_{1},\ldots,T_{m}$, where $T_{i}=T(\mathbf{X}_{n}^{(i)})$.
  3. 3.
    Apply a private mean estimator to the set of subsampled estimates $T_{1},\ldots,T_{m}$ to generate a private estimate of $T(\mathbf{X}_{n})$.
As mentioned previously, the private aggregator in subsample-and-aggregate algorithm is typically either the traditional private clipped mean, or the private widened winsorized mean applied coordinate-wise [29, 33]. If poor input bounds are provided for the coordinates of $T(\mathbf{X}_{n})$, this approach introduces a high level of noise to the private estimate. Moreover, we demonstrate that these algorithms can be unstable, or have large errors, when applied to small sample sizes, when the dimension is not very large. Critically, in the context of subsample-and-aggregate, we are often computing the mean of a small sample, given that we must first split the data into groups large enough for $T_{1},\ldots,T_{m}$ to be stable. This motivates our proposal of $\tilde{\mu}$ or $\tilde{\mu}^{\prime}$ applied coordinatewise as the aggregator. We demonstrate this observations below, via a real-world application to linear mixed modelling.
First, we state a corollary of Theorem 3.1, which yields the finite sample accuracy of the private estimate generated from subsample-and-aggregate. Assume that $\mathbf{X}_{n}^{\prime}$ is an adversarially corrupted version of $\mathbf{X}_{n}$, where at most $\lfloor n\eta\rfloor$ observations are corrupted. Let $\mathrm{Bias}_{n,\eta}(T)=\mathbb{E}(T(\mathbf{X}_{n}^{\prime}))-\theta$ be the bias of the estimator $T$. Next, given $T\colon\mathcal{D}_{n}\to\mathbb{R}$, let $\tilde{T}$ and $\tilde{T}^{\prime}$ be the subsample-and-aggregate estimates generated when $\tilde{\mu}$ and $\tilde{\mu}^{\prime}$ are used for the aggregation step, respectively. Let $\mathbf{X}_{n}^{\prime(1)},\ldots,\mathbf{X}_{n}^{\prime(m)}$ be a uniformly random set of disjoint subsets of size $k=\lfloor m/n\rfloor$ chosen from $\mathbf{X}_{n}^{\prime}$, and let $G_{k}$ be the distribution of $T(\mathbf{X}_{n}^{\prime(1)})$.
Corollary 5.1.
Given $T\colon\mathcal{D}_{n}\to\mathbb{R}$ and $1<m<n$, suppose that $T(\mathbf{X}_{n}^{\prime(1)})$ has variance $\sigma_{T,k}^{2}<\infty$, $\ell<\inf_{i\in[m]}T(\mathbf{X}_{n}^{\prime(i)})<\sup_{i\in[m]}T(\mathbf{X}_{n}^{\prime(i)})<u$ and that for all $\eta\geq 0$, $14e^{-n}<\delta<$ 1, $\varepsilon_{1},\varepsilon_{2}>112/3$, $\rho_{1},\rho_{2}>112/3$ and $\varepsilon_{3},\rho_{3}>0$, if $\beta,u,\ell\in\mathbb{R}$ are such that Condition 3 holds for $F=G_{k}$, then, with probability at least $1-\delta,$
$|\tilde{T}(\mathbf{X}_{n}^{\prime})-T(F)|\lesssim|\mathrm{Bias}_{n,\eta}(T)|+\sigma_{T,k}\left(\sqrt{\eta}\vee\sqrt{\frac{(\log(4\beta(u-\ell)(\beta-1))\vee 4)/\delta)}{m}}\vee\frac{1}{\sqrt{m}\varepsilon_{3}}\right).$ (5.1)
and
$|\tilde{T}(\mathbf{X}_{n}^{\prime})^{\prime}-T(F)|\lesssim|\mathrm{Bias}_{n,\eta}(T)|+\sigma_{T,k}\left(\sqrt{\eta}\vee\sqrt{\frac{(\log(4\beta(u-\ell)(\beta-1))\vee 4)/\delta)}{m}}\vee\frac{1}{\sqrt{m\rho_{3}}}\right).$ (5.2)
Proof.
First, the triangle inequality implies $|\tilde{T}(\mathbf{X}_{n}^{\prime})^{\prime}-T(F)|<|\mathrm{Bias}_{n,\eta}(T)|+|\tilde{T}(\mathbf{X}_{n}^{\prime})^{\prime}-\mathbb{E}(T(\mathbf{X}_{n}^{\prime(1)}))|$. Next, given that the conditions are satisfied by assumption, we apply Theorem 3.1. ∎
Note that we don’t necessarily need the restriction $\ell<\inf_{i\in[m]}T(\mathbf{X}_{n}^{\prime(i)})<\sup_{i\in[m]}T(\mathbf{X}_{n}^{\prime(i)})<u$, as this is stronger than what is required by Condition 2, especially in light of Remark 3.4. However, this condition makes the theorem easier to read and, since the bounds can be loose in practice, this restriction is not a practical concern.
Table 3: Empirical mean squared error between the non-private mixed model parameter estimates and each of the mixed model parameter estimates, produced by pairing subsample-and-aggregate algorithms with different private mean estimation algorithms, for different group sizes $k$ at $\rho=1$.
$k$ $\tilde{\mu}^{\prime}$ Clipped mean PRIME [25] Huber [35] Coinpress [23]
10 0.451 2.500 2290.289 2.929 2.341
20 0.214 4.744 2291.248 3.099 19.857
30 0.071 6.604 2293.800 3.135 69.541
40 0.057 9.186 2296.418 3.210 169.355
50 0.048 11.349 2295.262 3.221 417.814
60 0.056 14.193 2295.204 3.221 808.827
70 0.070 15.369 2301.360 3.227 1332.139
80 0.086 18.046 2308.593 3.228 1861.837
90 0.104 19.536 2311.090 3.230 3076.970
100 0.142 21.159 2310.142 3.228 4399.515
$\rho$
$\infty$ $10$ $1$
Winter 2.48 2.95 4.80
Time -12.18 -11.79 -11.95
M.L. 5.13 5.28 4.09
$\sigma^{2}_{B}$ 0.09 0.09 0.06
$\sigma^{2}_{W}$ 0.04 0.04 0.04
Figure 2: Left: Actual mean values, non-private estimated mean values and private estimated mean values of K10 scores of healthcare workers throughout the course of the COVID-19 pandemic. Right: Estimated percentage change in K10 scores for a one unit increase in each regressor, along with the estimated between ($\sigma^{2}_{B}$) and within ($\sigma^{2}_{W}$) subject variances, for different values of $\rho$, including the non-private estimates ($\rho=\infty$). Here, M.L. refers to mitigation level.
Corollary 5.1 yields two important insights. First, the optimal choice of $k$ depends on the bias of $T$, and, second, under contamination, the rate of convergence of $\tilde{T}$ depends on the robustness of $T$. Suppose that $\eta=0$, for expositional clarity on the first point. Taking the privacy parameters, $\ell$, $u$, and $\beta$ to be fixed, the rate of convergence is determined by two competing parameters, $|\mathrm{Bias}_{n,\eta}(T)|$ and $\sigma_{T,k}/m^{1/2}$. Provided that $\sigma_{T,k}=O(k^{-1/2})$, then $\sigma_{T,k}/m^{1/2}=O(n^{-1/2})$, which is the minimax optimal rate of convergence for many univariate problems.
To elaborate, by choosing a statistic $T$ that has a variance $O(k^{-1/2})$, the error of subsample-and-aggregate is bounded by $|\mathrm{Bias}_{n,\eta}(T)|+O(n^{-1/2})$. It remains then to select $k$ large enough so that $|\mathrm{Bias}_{n,\eta}(T)|=O(n^{-1/2})$. If $T$ is unbiased, then any choice of $k$ is sufficient for a convergence rate of $O(n^{-1/2})$. If $T$ is biased, then $k$ has to be large enough so that the bias is negligible. For instance, supposing that $F=\mathcal{N}(0,\sigma^{2})$, and taking $T$ to be the maximum likelihood estimator of the variance of the normal distribution, $|\mathrm{Bias}_{n,\eta}(T)|=\sigma^{2}/k$ and $\sigma_{T,k}\lesssim 1/\sqrt{k}$. In this case, taking $k\gtrsim\sqrt{n}$ results in a convergence rate of $O(n^{-1/2})$, which is the optimal convergence rate in the nonprivate setting.
For the second point, aside from the $\sqrt{\eta}$ term resulting from the mean estimation, note that the level of contamination is reflected in the contamination bias of $T$. Namely, if the sample is contaminated, in general, $|\mathrm{Bias}_{n,\eta}(T)|\leq|\mathbb{E}(T_{i}(\mathbf{x}_{k}^{\prime}))-\theta|+\sup_{\mathbf{x}_{k}}|\mathbb{E}(T_{i}(\mathbf{x}_{k}^{\prime}))-\mathbb{E}(T_{i}(\mathbf{x}_{k}))|$. This is the sum of the bias of $T$ in the uncontaminated setting and the contamination bias of $T$ for the dataset $\mathbf{x}_{k}^{\prime}$. If contamination is suspected, it is necessary to select a statistic $T$ that is “robust with high probability,” in the sense that for most datasets, $T$ is not easily contaminated. This is different from the stricter requirement of small global sensitivity, where it is required that $\sup_{\mathbf{x}_{k},\mathbf{x}_{k}^{\prime}}|T_{i}(\mathbf{x}_{k}^{\prime})-T_{i}(\mathbf{x}_{k})|$ is bounded. For example, the univariate median has infinite global sensitivity, but will generally be robust with high probability [14, 2, 32].
We now turn to an application of the improved subsample-and-aggregate. We consider the data of roughly 2500 healthcare workers, whose K10 combined depression and anxiety scores were tracked throughout the course of the COVID-19 pandemic [19]. One of the aims of the study was to estimate the effects of time, season, and periods of high mitigation on K10 scores. Given that this is longitudinal data, the authors applied a linear mixed effects model, specifically a random intercept model, to the log-transformed K10 scores. We estimate the same model privately, using the proposed improved subsample-and-aggregate. Specifically, we estimate the coefficients of time, winter, and mitigation level, as well as the intercept, the between subject variance, and the within subject variance. We further compare the output to the original subsample-and-aggregate, and subsample-and-aggregate algorithms where the aggregate step uses one of several state-of-the-art private multivariate mean estimation methods: PRIME [25], the private Huber estimator [35], or the Coinpress estimator [23]. Table 3 shows the empirical mean squared error between each of the private estimates and the non-private estimates, for different group sizes $k$ with $\rho=10$. Here, the provided input bounds for all algorithms were $\ell=-40\sqrt{6}=-u$. The full parameter settings are provided in Appendix B. In these results $\tilde{\mu}^{\prime}$ outperforms the original subsample-and-aggregate, and the multivariate methods. In addition, several multivariate methods fail for this application. One explanation is that, to achieve minimax optimality in the preceding settings, one must take into account the dependence structure of the data. For instance, many of the existing multivariate estimators privately estimate the covariance matrix [4, 6, 12]. While these estimators generally perform very well, at small sample sizes this estimate of the covariance matrix can become unstable, leading to unstable mean estimates.
Next, Figure 2 displays the estimated mean K10 scores using the private and non-private estimates. The rolling, four-week means of the raw data are also included. The private estimates are very close to the non-private ones, and all estimated means closely match the means of the raw data. Also, the table in Figure 2 compares the private parameter estimates generated from the improved subsample-and-aggregate, with the non-private estimates. Even at high levels of privacy, (i.e., $\rho=1$), the estimates align closely with the non-private estimates. The direction and magnitude of all effects are very similar, and the fitted mean functions align well with the data. We can conclude that, if the researchers had taken a differentially private approach, the conclusions of this investigation would not have changed materially. In general, we stress that subsample-and-aggregate is a powerful algorithm. Complex algorithms and models can be made private easily, and with minimal additional programming. By removing the need for tight input bounds on the data, or the need for large sample sizes, we unlock private versions of virtually any algorithm, without the need to design a new procedure for every different algorithm we may wish to privatize.
Appendix A Complete simulation results
This section contains tables which summarize all of the simulation results.
A.1 Results comparing subsample and aggregate in PDP
The following table contains the mean absolute error of the coefficients in the mixed modelling example in Section 5, generated via the subsample-and-aggregate algorithm with $\tilde{\mu}$ and the widened winsorized mean of Smith, [33] as the aggregators. We took $\epsilon=1$ as the privacy budget.
$k$ $\tilde{\mu}$ Widened Winsorized Mean [33]
10 0.453 2219.514
20 0.204 3678.827
30 0.145 3779.974
40 12.275 4339.895
50 85.240 3883.693
60 253.367 5157.979
70 327.143 5251.000
80 444.972 5968.068
90 530.348 7485.897
100 541.654 6908.987
A.2 Results comparing estimators
Below are the empirical mean squared errors of each private estimator, for each of the simulation scenarios.
$n=50$
Distribution $\rho$ $\tilde{\mu}$ $\tilde{\mu},\ \eta=0.3$ Bun, $t=0.01$ Bun, $t=1$ Duchi Kamath
Gaussian 1 0.0298 0.0400 3.8631 14.5905 24.4796 nan
Gaussian Mix. 1 0.6456 10.3692 4.1014 15.1445 25.2465 nan
Skewed 1 0.0398 0.0927 3.8373 15.0478 24.5744 nan
Heavy Tails 1 0.0564 0.0575 3.8699 14.9780 24.5289 nan
Cont. Gaussian 1 2.1576 0.1883 7.1301 18.7811 28.4708 nan
Gaussian 10 0.0208 0.0260 0.3724 0.0380 7.7342 nan
Gaussian Mix. 10 0.4944 1.4166 0.8024 0.5347 8.3742 nan
Skewed 10 0.0252 0.0738 0.3681 0.0407 7.7891 nan
Heavy Tails 10 0.0434 0.0399 0.3888 0.0614 7.7747 nan
Cont. Gaussian 10 3.9314 0.1712 3.8261 3.5541 11.7172 nan
Gaussian 100 0.0202 0.0243 0.0547 0.0207 2.4479 nan
Gaussian Mix. 100 0.4959 0.7021 0.5420 0.5309 3.0167 nan
Skewed 100 0.0235 0.0712 0.0567 0.0251 2.4805 nan
Heavy Tails 100 0.0456 0.0378 0.0749 0.0428 2.4836 nan
Cont. Gaussian 100 3.9856 0.1735 3.5638 3.5514 6.4264 nan
$n=100$
Distribution $\rho$ $\tilde{\mu}$ $\tilde{\mu},\ \eta=0.3$ Bun, $t=0.01$ Bun, $t=1$ Duchi Kamath
Gaussian 1 0.0117 0.0169 0.9439 0.9646 12.2418 53.8875
Gaussian Mix. 1 0.3011 1.8062 1.1915 1.5190 12.7386 60.0149
Skewed 1 0.0161 0.0688 0.9246 0.9784 12.2865 53.4350
Heavy Tails 1 0.0207 0.0214 0.9200 1.0200 12.4220 54.7056
Cont. Gaussian 1 3.7994 0.1466 4.1855 4.5279 16.2082 54.5297
Gaussian 10 0.0105 0.0127 0.0980 0.0118 3.8683 4.5231
Gaussian Mix. 10 0.2887 0.3923 0.3845 0.3117 4.2685 6.3717
Skewed 10 0.0117 0.0637 0.0949 0.0155 3.8938 4.4541
Heavy Tails 10 0.0188 0.0166 0.0952 0.0183 3.9757 4.6161
Cont. Gaussian 10 3.9517 0.1536 3.4365 3.3839 7.8290 7.5023
Gaussian 100 0.0104 0.0122 0.0197 0.0106 1.2248 0.5010
Gaussian Mix. 100 0.2868 0.3432 0.3174 0.3128 1.5706 2.1273
Skewed 100 0.0111 0.0619 0.0216 0.0147 1.2394 0.4924
Heavy Tails 100 0.0207 0.0155 0.0235 0.0173 1.2912 0.5435
Cont. Gaussian 100 3.9575 0.1547 3.3870 3.3893 5.1822 3.1129
$n=500$
Distribution $\rho$ $\tilde{\mu}$ $\tilde{\mu},\ \eta=0.3$ Bun, $t=0.01$ Bun, $t=1$ Duchi Kamath
Gaussian 1 0.0026 0.0029 0.0270 0.0172 2.4723 1.4319
Gaussian Mix. 1 0.0468 0.0550 0.0720 0.0743 2.5221 1.5452
Skewed 1 0.0026 0.0570 0.0314 0.0213 2.4553 1.4297
Heavy Tails 1 0.0045 0.0036 0.0289 0.0213 2.4697 1.4758
Cont. Gaussian 1 3.9835 0.1522 3.3974 3.4539 6.4617 5.0107
Gaussian 10 0.0025 0.0028 0.0047 0.0025 0.7873 0.1532
Gaussian Mix. 10 0.0464 0.0544 0.0520 0.0507 0.8346 0.2774
Skewed 10 0.0023 0.0568 0.0093 0.0072 0.7775 0.1475
Heavy Tails 10 0.0045 0.0034 0.0062 0.0038 0.7870 0.1545
Cont. Gaussian 10 3.9777 0.1521 3.3958 3.4032 4.7742 4.0160
Gaussian 100 0.0025 0.0028 0.0027 0.0025 0.2528 0.0211
Gaussian Mix. 100 0.0465 0.0545 0.0506 0.0507 0.2988 0.1641
Skewed 100 0.0023 0.0565 0.0074 0.0072 0.2470 0.0204
Heavy Tails 100 0.0044 0.0034 0.0040 0.0038 0.2538 0.0244
Cont. Gaussian 100 3.9778 0.1518 3.4000 3.4026 4.2382 3.9773
$n=1000$
Distribution $\rho$ $\tilde{\mu}$ $\tilde{\mu},\ \eta=0.3$ Bun, $t=0.01$ Bun, $t=1$ Duchi Kamath
Gaussian 1 0.0013 0.0015 0.0054 0.0050 1.2366 0.3324
Gaussian Mix. 1 0.0258 0.0317 0.0319 0.0365 1.2605 0.4570
Skewed 1 0.0015 0.0586 0.0107 0.0097 1.2295 0.3388
Heavy Tails 1 0.0021 0.0016 0.0064 0.0053 1.2385 0.3351
Cont. Gaussian 1 3.9906 0.1543 3.4036 3.4289 5.2380 4.2162
Gaussian 10 0.0013 0.0015 0.0017 0.0013 0.3939 0.0368
Gaussian Mix. 10 0.0257 0.0308 0.0286 0.0285 0.4181 0.1191
Skewed 10 0.0014 0.0587 0.0071 0.0068 0.3898 0.0362
Heavy Tails 10 0.0022 0.0015 0.0024 0.0018 0.3957 0.0372
Cont. Gaussian 10 3.9915 0.1539 3.4104 3.4146 4.3935 3.9491
Gaussian 100 0.0013 0.0015 0.0013 0.0013 0.1266 0.0064
Gaussian Mix. 100 0.0257 0.0310 0.0284 0.0284 0.1509 0.0836
Skewed 100 0.0014 0.0587 0.0068 0.0068 0.1241 0.0059
Heavy Tails 100 0.0022 0.0015 0.0019 0.0018 0.1283 0.0089
Cont. Gaussian 100 3.9910 0.1548 3.4133 3.4146 4.1251 3.9003
A.3 Results comparing the parameters $\rho$ and $C$
Below are the empirical mean squared errors of $\tilde{\mu}^{\prime}$, for different values of $\eta$ and $C$.
$n=50$
Distribution $\rho$ $C$ $\eta$=0.0 $\eta$=0.05 $\eta$=0.15 $\eta$=0.3
Gaussian 1 5 0.0288 0.0307 0.0339 0.0415
Gaussian 1 10 0.0288 0.0307 0.0339 0.0415
Gaussian 1 100 0.0288 0.0307 0.0339 0.0415
Gaussian Mix. 1 5 0.5326 0.7573 2.2593 10.1633
Gaussian Mix. 1 10 0.5326 0.7573 2.2593 10.1633
Gaussian Mix. 1 100 0.5326 0.7573 2.2593 10.1633
Skewed 1 5 0.0387 0.0415 0.0640 0.0932
Skewed 1 10 0.0387 0.0415 0.0640 0.0932
Skewed 1 100 0.0387 0.0415 0.0640 0.0932
Heavy Tails 1 5 0.0514 0.0583 0.0518 0.0572
Heavy Tails 1 10 0.0514 0.0583 0.0518 0.0572
Heavy Tails 1 100 0.0514 0.0583 0.0518 0.0572
Cont. Gaussian 1 5 2.0557 1.6832 0.5690 0.2016
Cont. Gaussian 1 10 2.0557 1.6832 0.5690 0.2016
Cont. Gaussian 1 100 2.0557 1.6832 0.5690 0.2016
Gaussian 10 5 0.0236 0.0230 0.0244 0.0276
Gaussian 10 10 0.0236 0.0230 0.0244 0.0276
Gaussian 10 100 0.0236 0.0230 0.0244 0.0276
Gaussian Mix. 10 5 0.5315 0.5175 0.5458 1.5674
Gaussian Mix. 10 10 0.5315 0.5175 0.5458 1.5674
Gaussian Mix. 10 100 0.5315 0.5438 0.5458 1.5674
Skewed 10 5 0.0228 0.0242 0.0398 0.0773
Skewed 10 10 0.0228 0.0242 0.0398 0.0773
Skewed 10 100 0.0228 0.0242 0.0398 0.0773
Heavy Tails 10 5 0.0521 0.0467 0.0384 0.0335
Heavy Tails 10 10 0.0521 0.0467 0.0384 0.0335
Heavy Tails 10 100 0.0521 0.0467 0.0384 0.0335
Cont. Gaussian 10 5 3.9661 3.9408 1.7375 0.1739
Cont. Gaussian 10 10 3.9661 3.9408 1.7375 0.1739
Cont. Gaussian 10 100 3.9661 3.9408 1.7375 0.1739
Gaussian 100 5 0.0224 0.0226 0.0235 0.0261
Gaussian 100 10 0.0224 0.0226 0.0235 0.0261
Gaussian 100 100 0.0224 0.0226 0.0235 0.0261
Gaussian Mix. 100 5 0.5161 0.5178 0.5279 0.7180
Gaussian Mix. 100 10 0.5161 0.5178 0.5279 0.7180
Gaussian Mix. 100 100 0.5161 0.5178 0.5279 0.7180
Skewed 100 5 0.0201 0.0216 0.0367 0.0749
Skewed 100 10 0.0201 0.0216 0.0367 0.0749
Skewed 100 100 0.0201 0.0216 0.0367 0.0749
Heavy Tails 100 5 0.0572 0.0494 0.0388 0.0315
Heavy Tails 100 10 0.0572 0.0494 0.0388 0.0315
Heavy Tails 100 100 0.0572 0.0494 0.0388 0.0315
Cont. Gaussian 100 5 3.9665 3.9475 3.6710 0.1727
Cont. Gaussian 100 10 3.9665 3.9475 3.6710 0.1727
Cont. Gaussian 100 100 3.9665 3.9475 3.6710 0.1727
$n=100$
Distribution $\rho$ $C$ $\eta$=0.0 $\eta$=0.05 $\eta$=0.15 $\eta$=0.3
Gaussian 1 5 0.0124 0.0123 0.0146 0.0170
Gaussian 1 10 0.0124 0.0123 0.0146 0.0170
Gaussian 1 100 0.0124 0.0123 0.0146 0.0170
Gaussian Mix. 1 5 0.2685 0.2701 0.2761 1.3378
Gaussian Mix. 1 10 0.2685 0.2701 0.2761 1.3378
Gaussian Mix. 1 100 0.2685 0.2701 0.2761 1.3378
Skewed 1 5 0.0156 0.0191 0.0359 0.0732
Skewed 1 10 0.0156 0.0191 0.0359 0.0732
Skewed 1 100 0.0156 0.0191 0.0359 0.0732
Heavy Tails 1 5 0.0226 0.0195 0.0180 0.0175
Heavy Tails 1 10 0.0226 0.0195 0.0180 0.0175
Heavy Tails 1 100 0.0226 0.0195 0.0180 0.0175
Cont. Gaussian 1 5 3.8225 3.5852 1.0125 0.1704
Cont. Gaussian 1 10 3.8225 3.5852 1.0125 0.1704
Cont. Gaussian 1 100 3.8225 3.5852 1.0125 0.1704
Gaussian 10 5 0.0112 0.0116 0.0121 0.0138
Gaussian 10 10 0.0112 0.0116 0.0121 0.0138
Gaussian 10 100 0.0112 0.0116 0.0121 0.0138
Gaussian Mix. 10 5 0.2576 0.2614 0.2688 0.3244
Gaussian Mix. 10 10 0.2576 0.2614 0.2688 0.3244
Gaussian Mix. 10 100 0.2576 0.2614 0.2688 0.3244
Skewed 10 5 0.0117 0.0137 0.0292 0.0682
Skewed 10 10 0.0117 0.0137 0.0292 0.0682
Skewed 10 100 0.0117 0.0137 0.0292 0.0682
Heavy Tails 10 5 0.0203 0.0187 0.0151 0.0144
Heavy Tails 10 10 0.0203 0.0187 0.0151 0.0144
Heavy Tails 10 100 0.0203 0.0187 0.0151 0.0144
Cont. Gaussian 10 5 3.9825 3.9533 3.3835 0.1651
Cont. Gaussian 10 10 3.9825 3.9533 3.3835 0.1651
Cont. Gaussian 10 100 3.9825 3.9533 3.3835 0.1651
Gaussian 100 5 0.0110 0.0112 0.0120 0.0133
Gaussian 100 10 0.0110 0.0112 0.0120 0.0133
Gaussian 100 100 0.0110 0.0112 0.0120 0.0133
Gaussian Mix. 100 5 0.2610 0.2608 0.2709 0.3055
Gaussian Mix. 100 10 0.2610 0.2608 0.2709 0.3055
Gaussian Mix. 100 100 0.2610 0.2608 0.2709 0.3055
Skewed 100 5 0.0111 0.0126 0.0282 0.0680
Skewed 100 10 0.0111 0.0126 0.0282 0.0680
Skewed 100 100 0.0111 0.0126 0.0282 0.0680
Heavy Tails 100 5 0.0216 0.0192 0.0157 0.0133
Heavy Tails 100 10 0.0216 0.0192 0.0157 0.0133
Heavy Tails 100 100 0.0216 0.0192 0.0157 0.0133
Cont. Gaussian 100 5 3.9827 3.9665 3.6897 0.1642
Cont. Gaussian 100 10 3.9827 3.9743 3.6897 0.1642
Cont. Gaussian 100 100 3.9827 3.9665 3.6897 0.1642
$n=500$
Distribution $\rho$ $C$ $\eta$=0.0 $\eta$=0.05 $\eta$=0.15 $\eta$=0.3
Gaussian 1 5 0.0020 0.0019 0.0020 0.0024
Gaussian 1 10 0.0019 0.0019 0.0020 0.0024
Gaussian 1 100 0.0020 0.0019 0.0020 0.0024
Gaussian Mix. 1 5 0.0507 0.0512 0.0536 0.0608
Gaussian Mix. 1 10 0.0514 0.0512 0.0536 0.0608
Gaussian Mix. 1 100 0.0523 0.0512 0.0536 0.0608
Skewed 1 5 0.0026 0.0046 0.0196 0.0575
Skewed 1 10 0.0028 0.0046 0.0196 0.0575
Skewed 1 100 0.0029 0.0046 0.0196 0.0575
Heavy Tails 1 5 0.0044 0.0036 0.0033 0.0032
Heavy Tails 1 10 0.0042 0.0036 0.0033 0.0032
Heavy Tails 1 100 0.0044 0.0036 0.0033 0.0032
Cont. Gaussian 1 5 3.9931 3.9535 3.6542 0.1539
Cont. Gaussian 1 10 3.9847 3.9535 3.6542 0.1539
Cont. Gaussian 1 100 3.9814 3.9535 3.6542 0.1539
Gaussian 10 5 0.0018 0.0019 0.0020 0.0022
Gaussian 10 10 0.0019 0.0019 0.0020 0.0022
Gaussian 10 100 0.0019 0.0019 0.0020 0.0022
Gaussian Mix. 10 5 0.0511 0.0518 0.0532 0.0603
Gaussian Mix. 10 10 0.0510 0.0518 0.0532 0.0603
Gaussian Mix. 10 100 0.0510 0.0518 0.0532 0.0603
Skewed 10 5 0.0024 0.0041 0.0193 0.0576
Skewed 10 10 0.0025 0.0041 0.0193 0.0576
Skewed 10 100 0.0027 0.0041 0.0193 0.0576
Heavy Tails 10 5 0.0043 0.0036 0.0031 0.0028
Heavy Tails 10 10 0.0044 0.0036 0.0031 0.0028
Heavy Tails 10 100 0.0042 0.0036 0.0031 0.0028
Cont. Gaussian 10 5 3.9915 3.9627 3.6861 0.1529
Cont. Gaussian 10 10 3.9874 3.9627 3.6861 0.1529
Cont. Gaussian 10 100 3.9849 3.9627 3.6861 0.1529
Gaussian 100 5 0.0018 0.0019 0.0019 0.0022
Gaussian 100 10 0.0019 0.0019 0.0019 0.0022
Gaussian 100 100 0.0019 0.0019 0.0019 0.0022
Gaussian Mix. 100 5 0.0513 0.0515 0.0535 0.0604
Gaussian Mix. 100 10 0.0511 0.0515 0.0535 0.0604
Gaussian Mix. 100 100 0.0512 0.0515 0.0535 0.0604
Skewed 100 5 0.0024 0.0040 0.0192 0.0580
Skewed 100 10 0.0025 0.0040 0.0192 0.0580
Skewed 100 100 0.0026 0.0040 0.0192 0.0580
Heavy Tails 100 5 0.0043 0.0037 0.0030 0.0028
Heavy Tails 100 10 0.0042 0.0037 0.0030 0.0028
Heavy Tails 100 100 0.0041 0.0037 0.0030 0.0028
Cont. Gaussian 100 5 3.9931 3.9621 3.6867 0.1533
Cont. Gaussian 100 10 3.9875 3.9621 3.6867 0.1533
Cont. Gaussian 100 100 3.9838 3.9621 3.6867 0.1533
$n=1000$
Distribution $\rho$ $C$ $\eta$=0.0 $\eta$=0.05 $\eta$=0.15 $\eta$=0.3
Gaussian 1 5 0.0011 0.0011 0.0011 0.0013
Gaussian 1 10 0.0011 0.0011 0.0011 0.0013
Gaussian 1 100 0.0010 0.0011 0.0011 0.0013
Gaussian Mix. 1 5 0.0258 0.0257 0.0260 0.0297
Gaussian Mix. 1 10 0.0254 0.0257 0.0260 0.0297
Gaussian Mix. 1 100 0.0251 0.0257 0.0260 0.0297
Skewed 1 5 0.0012 0.0034 0.0196 0.0588
Skewed 1 10 0.0012 0.0034 0.0196 0.0588
Skewed 1 100 0.0018 0.0034 0.0196 0.0588
Heavy Tails 1 5 0.0028 0.0024 0.0022 0.0018
Heavy Tails 1 10 0.0027 0.0024 0.0022 0.0018
Heavy Tails 1 100 0.0026 0.0024 0.0022 0.0018
Cont. Gaussian 1 5 3.9903 3.9629 3.6685 0.1509
Cont. Gaussian 1 10 3.9889 3.9629 3.6685 0.1509
Cont. Gaussian 1 100 3.9819 3.9629 3.6685 0.1509
Gaussian 10 5 0.0010 0.0010 0.0011 0.0012
Gaussian 10 10 0.0010 0.0010 0.0011 0.0012
Gaussian 10 100 0.0010 0.0010 0.0011 0.0012
Gaussian Mix. 10 5 0.0254 0.0254 0.0261 0.0295
Gaussian Mix. 10 10 0.0254 0.0254 0.0261 0.0295
Gaussian Mix. 10 100 0.0253 0.0254 0.0261 0.0295
Skewed 10 5 0.0011 0.0032 0.0195 0.0592
Skewed 10 10 0.0012 0.0032 0.0195 0.0592
Skewed 10 100 0.0016 0.0032 0.0195 0.0592
Heavy Tails 10 5 0.0027 0.0024 0.0021 0.0017
Heavy Tails 10 10 0.0026 0.0024 0.0021 0.0017
Heavy Tails 10 100 0.0025 0.0024 0.0021 0.0017
Cont. Gaussian 10 5 3.9917 3.9624 3.6782 0.1511
Cont. Gaussian 10 10 3.9915 3.9624 3.6782 0.1511
Cont. Gaussian 10 100 3.9823 3.9624 3.6782 0.1511
Gaussian 100 5 0.0010 0.0010 0.0011 0.0012
Gaussian 100 10 0.0010 0.0010 0.0011 0.0012
Gaussian 100 100 0.0010 0.0010 0.0011 0.0012
Gaussian Mix. 100 5 0.0253 0.0254 0.0261 0.0293
Gaussian Mix. 100 10 0.0254 0.0254 0.0261 0.0293
Gaussian Mix. 100 100 0.0254 0.0254 0.0261 0.0293
Skewed 100 5 0.0011 0.0033 0.0196 0.0594
Skewed 100 10 0.0012 0.0033 0.0196 0.0594
Skewed 100 100 0.0016 0.0033 0.0196 0.0594
Heavy Tails 100 5 0.0028 0.0024 0.0021 0.0017
Heavy Tails 100 10 0.0026 0.0024 0.0021 0.0017
Heavy Tails 100 100 0.0025 0.0024 0.0021 0.0017
Cont. Gaussian 100 5 3.9925 3.9617 3.6799 0.1512
Cont. Gaussian 100 10 3.9909 3.9617 3.6799 0.1512
Cont. Gaussian 100 100 3.9828 3.9617 3.6799 0.1512
A.4 Results concerning the privacy budget distribution
Below are the empirical mean squared errors of $\tilde{\mu}^{\prime}$, for different distributions of $\rho$, across the quantile estimation step and the mean estimation step.
$n=50$
Distribution $\rho$ Quantile favored Equal Mean favored
Gaussian 1 0.0358 0.0332 0.0334
Gaussian Mix. 1 0.7170 0.6965 0.7507
Skewed 1 0.0308 0.0293 0.0452
Heavy Tails 1 0.0681 0.0496 0.0522
Cont. Gaussian 1 3.4910 3.0430 1.9391
Gaussian 10 0.0190 0.0183 0.0184
Gaussian Mix. 10 0.5764 0.5750 0.5784
Skewed 10 0.0191 0.0191 0.0197
Heavy Tails 10 0.0432 0.0406 0.0389
Cont. Gaussian 10 3.9924 3.9523 3.9383
Gaussian 100 0.0183 0.0181 0.0183
Gaussian Mix. 100 0.5606 0.5605 0.5624
Skewed 100 0.0174 0.0173 0.0175
Heavy Tails 100 0.0398 0.0396 0.0405
Cont. Gaussian 100 3.9660 3.9667 3.9638
$n=100$
Distribution $\rho$ Quantile favored Equal Mean favored
Gaussian 1 0.0123 0.0121 0.0130
Gaussian Mix. 1 0.2806 0.2693 0.2857
Skewed 1 0.0119 0.0128 0.0133
Heavy Tails 1 0.0284 0.0250 0.0265
Cont. Gaussian 1 4.0326 3.9716 3.8126
Gaussian 10 0.0109 0.0108 0.0107
Gaussian Mix. 10 0.2533 0.2515 0.2522
Skewed 10 0.0090 0.0088 0.0091
Heavy Tails 10 0.0236 0.0235 0.0228
Cont. Gaussian 10 3.9815 3.9862 3.9770
Gaussian 100 0.0105 0.0104 0.0105
Gaussian Mix. 100 0.2520 0.2513 0.2512
Skewed 100 0.0087 0.0086 0.0087
Heavy Tails 100 0.0248 0.0247 0.0244
Cont. Gaussian 100 3.9752 3.9713 3.9729
$n=500$
Distribution $\rho$ Quantile favored Equal Mean favored
Gaussian 1 0.0024 0.0023 0.0023
Gaussian Mix. 1 0.0547 0.0549 0.0548
Skewed 1 0.0028 0.0027 0.0029
Heavy Tails 1 0.0060 0.0051 0.0059
Cont. Gaussian 1 4.0029 4.0041 4.0079
Gaussian 10 0.0021 0.0021 0.0021
Gaussian Mix. 10 0.0545 0.0544 0.0545
Skewed 10 0.0024 0.0024 0.0025
Heavy Tails 10 0.0053 0.0053 0.0053
Cont. Gaussian 10 3.9975 3.9990 3.9977
Gaussian 100 0.0021 0.0021 0.0021
Gaussian Mix. 100 0.0546 0.0546 0.0546
Skewed 100 0.0024 0.0024 0.0024
Heavy Tails 100 0.0056 0.0056 0.0055
Cont. Gaussian 100 3.9966 3.9969 3.9969
$n=1000$
Distribution $\rho$ Quantile favored Equal Mean favored
Gaussian 1 0.0011 0.0011 0.0011
Gaussian Mix. 1 0.0256 0.0260 0.0258
Skewed 1 0.0015 0.0015 0.0015
Heavy Tails 1 0.0027 0.0022 0.0021
Cont. Gaussian 1 3.9928 3.9931 4.0017
Gaussian 10 0.0011 0.0011 0.0011
Gaussian Mix. 10 0.0259 0.0259 0.0260
Skewed 10 0.0014 0.0014 0.0014
Heavy Tails 10 0.0021 0.0022 0.0021
Cont. Gaussian 10 3.9968 3.9975 3.9975
Gaussian 100 0.0011 0.0011 0.0011
Gaussian Mix. 100 0.0260 0.0260 0.0260
Skewed 100 0.0014 0.0014 0.0014
Heavy Tails 100 0.0021 0.0021 0.0021
Cont. Gaussian 100 3.9974 3.9975 3.9977
Appendix B Full description of subsample-and-aggregate parameter settings
Below are the full parameter settings used to create Table 3. Each scenario was ran 50 times.
  • Where required, the bound on the norm of the vector of estimates was set to $2\times 40\sqrt{d}=80\sqrt{6}$.
  • Where required, the lower bound on each coordinate was set to $\ell=-40\sqrt{6}$, and the upper bound was set to $u=40\sqrt{6}$.
  • For the estimator of Yu et al., [35]: $T$ was set to $\lfloor\log n\rfloor$ and other parameters were left at the default values, as per the code provided kindly from Yu et al., [35].
  • For $\tilde{\mu}^{\prime}$: $C=1$, $\eta=0$ and $\beta=1.001$.
  • For the estimator of Biswas et al., [4]: the upper bound on the covariance was set to $50\sqrt{d}=50\sqrt{6}$ and the starting center was set to the 0 vector. Also, $R=80\sqrt{6}$ and $\beta=0.1$. Three iterations were used in both the mean and covariance steps, with $\rho$ set to $1/3c,1/2c,1/c$ with $c=(1/2+1/3+1)/2$.
  • For the estimator of Liu et al., 2021a [25]: $\alpha=0.125$, $R=80\sqrt{6}$, and $\delta=0.01$ and other parameters were left at the default values, as per the code provided kindly from Liu et al., 2021a [25].
Appendix C Proof of Theorem 3.1
We now prove Theorem 3.1. Let $\tilde{X}=X-\mu$, $\xi_{q,X}$ be the (left continuous) $q$th quantile of the distribution of the random variable $X\sim F$, $\hat{\xi}_{q,X}$ be the (non-private) estimated $q$th quantile based on $X$, and $\tilde{\xi}_{q,X}$ the private estimate $q$th quantile based on a random sample of size $n$, whose observations have the same distribution as $X$.
Proof of Theorem 3.1.
We prove the result for pure differential privacy, where the result for $\tilde{\mu}^{\prime}$ follows by the same logic. Consider the following inequalities:
$\xi_{1-7\zeta/4}\leq\hat{\xi}_{1-3\zeta/2}\leq\xi_{1-5\zeta/4}\leq\xi_{1-3\zeta/4}\leq\hat{\xi}_{1-\zeta/2}\leq\xi_{1-\zeta/4}$ (C.1)
$\xi_{\zeta/4}\leq\hat{\xi}_{\zeta/2}\leq\xi_{3\zeta/4}\leq\xi_{5\zeta/4}\leq\hat{\xi}_{3\zeta/2}\leq\xi_{7\zeta/4}.$ (C.2)
Let the event $E$ be the event on which the above inequalities hold. The first step of the proof is to show that on $E$, we have that
$\hat{\xi}_{\zeta/2}\leq\tilde{\xi}_{\zeta,Y^{\prime}}\leq\hat{\xi}_{3\zeta/2}\qquad\text{and}\qquad\hat{\xi}_{1-3\zeta/2}\leq\tilde{\xi}_{1-\zeta,Y^{\prime}}\leq\hat{\xi}_{1-\zeta/2},$ (C.3)
holds with high probability. Then, we will show that $E$ occurs with high probability.
The inequality (C.3) is implied by $\zeta/2\leq F_{n}(\tilde{\xi}_{\zeta,Y^{\prime}})\leq\frac{3}{2}\zeta$ and $21-\frac{3}{2}\zeta\leq F_{n}(\tilde{\xi}_{1-\zeta,Y^{\prime}})\leq 1-\zeta/$ 2, where one may recall that $F_{n}(x)=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\leq x\right\}/n$ for any $x\in\mathbb{R}$. Note that these inequalities are implied by the statements “$F_{n}(\tilde{\xi}_{\zeta})$ is within $\zeta/2$ of $\zeta$”, and that “$F_{n}(\tilde{\xi}_{1-\zeta})$ is within $\zeta/2$ of $1-\zeta$”, respectively. That is, taking $q=\zeta$ and $t=\zeta/2$, we are concerned with the quantities $|q-F_{n}(\tilde{\xi}_{q})|<t$ and $|1-q-F_{n}(\tilde{\xi}_{1-q})|<t$. To prove that these inequalities hold with high probability, we can apply Lemma E.1, once taking $q=\zeta$ and $t=\zeta/2$ and again with $q=1-\zeta$ and $t=\zeta/2$.
In order to apply Lemma E.1, we must have then that three inequalities hold. Namely, $\zeta<1$, and
$\beta$ $\leq\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1},\text{and}$ (C.4)
$\beta$ $\leq\frac{\hat{\xi}_{1-\zeta/2}-\ell+1}{\hat{\xi}_{1-3\zeta/2}-\ell+1}.$ (C.5)
First, $\zeta<1$ holds by Condition 2. For (C.4), note that (C.2) directly implies
$\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1}$ $=1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\hat{\xi}_{3\zeta/2}+1}>1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\xi_{5\zeta/4}+1}>1+\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}>1+b_{n}.$
Similarly, for (C.5), (C.1) directly implies that
$\frac{\hat{\xi}_{1-\zeta/2}-\ell+1}{\hat{\xi}_{1-3\zeta/2}-\ell+1}$ $=1+\frac{\hat{\xi}_{1-\zeta/2}-\hat{\xi}_{1-3\zeta/2}}{\hat{\xi}_{1-3\zeta/2}-\ell+1}>1+\frac{\xi_{1-3\zeta/4}-\xi_{1-5\zeta/4}}{\xi_{1-5\zeta/4}-\ell+1}>1+b_{n}.$
Condition 3 directly says that $\beta<1+b_{n}$, giving (C.4) and (C.5). Thus, we can apply Lemma E.1. Applying Lemma E.1 gives
$\Pr\left(\left|\zeta-F_{n}(\tilde{\xi}_{\zeta})\right|>\zeta/2\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2),\quad\text{and}$
$\Pr\left(\left|1-\frac{\zeta}{2}-F_{n}(\tilde{\xi}_{1-\zeta/2})\right|>\frac{\zeta}{2}\right)$ $\leq\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2).$
As a result, on $E$, (C.3) fails to hold with probability at most,
$\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)+\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
$=\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)+\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+2\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
$\leq\left(\frac{\log(u-\xi_{5\zeta/4}+1)+\log(\xi_{1-5\zeta/4}-\ell+1)}{\log\beta}+2\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
$\leq 2\left(\frac{\log(u-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2).$
The final inequality follows applying (C.1) and (C.2) and Condition 2. Consider that, since $\varepsilon_{1}\wedge\varepsilon_{2}\geq 3/56$ and
$\zeta\leq\frac{112}{3n}\log\left(\frac{32}{\delta}\left[\frac{\beta(u-\ell)}{\beta-1}\vee 1\right]\right).$
Combining these results yields
$2\left(\frac{\log((u-\ell)+1)}{\log\beta}+1\right)\exp\left(-\frac{n\zeta(\varepsilon_{1}\wedge\varepsilon_{2})}{2}\right)$
$\leq 2\left(\frac{\log((u-\ell)+1)}{\log\beta}+1\right)\exp\left(-\frac{n\frac{112}{3n}\log\left(\frac{32}{\delta}\left[\frac{\beta(u-\ell)}{\beta-1}\vee 1\right]\right)(3/56)}{2}\right)$
$\leq 2\left(\frac{(u-\ell)\beta}{\beta-1}+1\right)\frac{\delta}{32}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}\leq\frac{\delta}{8}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}=\frac{\delta}{8}.$
Thus, combining the preceding statements yields that, on $E$, we have that (C.3) holds with probability at least
$1-2\left(\frac{\log((u-\ell)+1)}{\log\beta}+1\right)\exp\left(-\frac{n\zeta(\varepsilon_{1}\wedge\varepsilon_{2})}{2}\right)\geq 1-\frac{\delta}{8}.$
The next step is to show that $E$ occurs, i.e., the inequalities (C.1) and (C.2) hold, with high probability. Call the event on which the inequalities (C.3) hold $A_{1}$. The inequalities (C.1) and (C.2) when combined with (C.3) imply,
$\xi_{\zeta/4,}<\tilde{\xi}_{\zeta,Y^{\prime}}<\xi_{7\zeta/4}<\xi_{4\zeta}\qquad\text{and}\qquad\xi_{1-4\zeta}<\xi_{1-7\zeta/4}<\tilde{\xi}_{1-\zeta,Y^{\prime}}<\xi_{1-\zeta/4}.$
The next portion of the proof is similar to that of Theorem 1 of Lugosi and Mendelson, [28], however some steps are different, as that proof contains an error, see Appendix F. Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\})\text{ and }\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, [28]. To this end, observe that $\Pr\left(\tilde{X}\geq\xi_{1-7\zeta/4,\tilde{X}}\right)=7\zeta/4$. Recall that $\tilde{X}=X-\mu$. Define $W_{k-}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\leq\mu+\xi_{k,\tilde{X}}\right\}$ and $W_{k+}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\geq\mu+\xi_{k,\tilde{X}}\right\}$. Then, note that
$\Pr\left(W_{k\zeta-}\leq an\zeta\right)$ $=\Pr\left(W_{k\zeta-}-nk\zeta\leq n\zeta(a-k)\right)$
$=\Pr\left(nk-W_{k\zeta-}\geq n\zeta(k-a)\right)$
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(k-a)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(k-a)/3}\right)=\exp\left(-\frac{n\zeta(k-a)^{2}/2}{k(1-k\zeta)+(k-a)/3}\right).$
This holds so long as $k\geq a$, by Bernstein’s inequality. The same argument holds for $P(W_{(1-k\zeta)+}\leq an\zeta)$, giving the same upper bound. Effectively the same argument gives
$\Pr\left(W_{(1-k\zeta)-}\leq n(1-a^{\prime}\zeta)\right)$ $=\Pr\left(W_{(1-k\zeta)-}-n(1-k\zeta)\leq n(1-a^{\prime}\zeta-(1-k\zeta))\right)$
$=\Pr\left(n(1-k\zeta)-W_{(1-k\zeta)-}\geq n\zeta(a^{\prime}-k)\right)$
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(a^{\prime}-k)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(a^{\prime}-k)/3}\right)=\exp\left(-\frac{n\zeta(a^{\prime}-k)^{2}/2}{k(1-k\zeta)+(a^{\prime}-k)/3}\right).$
This holds so long as $k\leq a^{\prime}$. Symmetric arguments give the same bound for $P(W_{k\zeta+}\leq n(1-a^{\prime}\zeta))$.
Suppose $W_{1}=W_{(1-4\zeta)+}$, $W_{2}=W_{(1-\zeta/4)-}$, $W_{3}=W_{4\zeta-}$, and $W_{4}=W_{\zeta/4+}$. Then, using the above results,
$P(W_{1}\leq 17n\zeta/8)$ $\leq\exp\left(-\frac{n\zeta(4-17/8)^{2}/2}{4(1-4\zeta)+(4-17/8)/3}\right)=\exp\left(-\frac{225n\zeta}{592-2048\zeta}\right)$
$P(W_{2}\leq n(1-3\zeta/8))$ $\leq\exp\left(-\frac{n\zeta(3/8-1/4)^{2}/2}{1/4(1-1/4\zeta)+(3/8-1/4)/3}\right)=\exp\left(-\frac{3n\zeta}{112-24\zeta}\right)$
$P(W_{3}\leq 17n\zeta/8)$ $\leq\exp\left(-\frac{n\zeta(4-17/8)^{2}/2}{4(1-4\zeta)+(4-17/8)/3}\right)=\exp\left(-\frac{225n\zeta}{592-2048\zeta}\right)$
$P(W_{4}\leq n(1-3\zeta/8))$ $\leq\exp\left(-\frac{n\zeta(3/8-1/4)^{2}/2}{1/4(1-1/4\zeta)+(3/8-1/4)/3}\right)=\exp\left(-\frac{3n\zeta}{112-24\zeta}\right).$
Note that, for $40\leq\zeta\leq 1/$ 4,
$\exp\left(-\frac{225n\zeta}{592-2048\zeta}\right)\leq\exp\left(-\frac{3n\zeta}{112-24\zeta}\right).$
Thus, with probability at least $1-4\exp\left(-\frac{3n\zeta}{112-24\zeta}\right)\geq 1-4\exp\left(-\frac{3n\zeta}{112}\right)$, we have that
$A_{2}=\{W_{1}\geq 17n\zeta/8\}\cap\{W_{2}\geq n(1-3\zeta/8)\}\cap\{W_{3}\geq 17n\zeta/8\}\cap\{W_{4}\geq n(1-3\zeta/8)\}$
holds.
Recall that
$\zeta=16\eta+\frac{112}{3}\frac{\log(32(\beta(u-\ell)/(\beta-1)\vee 1)/\delta)}{n}\geq\frac{112}{3}\frac{\log(32/\delta)}{n}.$
Then, plugging this bound into the previously derived probability, we find that $A_{2}$ holds with probability at least $81-\delta/$ 8.
Now, using the fact that $\eta\leq\zeta/16$, on $A_{2}$, we have that the following inequalities hold:
$W_{1}\geq\frac{17}{8}\zeta n\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-4\zeta,\tilde{X}}\right\}\geq\frac{17}{8}\zeta n-2\eta n\geq\frac{17}{8}\zeta n-n\zeta/8\geq 2n\zeta,$
$W_{2}\geq(1-3\zeta/8)n\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-\zeta/4,\tilde{X}}\right\}\geq(1-3\zeta/8)n-2\eta n\geq n(1-\zeta/2).$
Similar arguments give $W_{3}\geq 2n\zeta$ and $W_{4}\geq n(1-\zeta/2)$, on $A_{2}$.
These inequalities state that, on $A_{2}$, the number of centered points in the corrupted sample above $\xi_{1-4\zeta,\tilde{X}}$ is at least $2n\zeta$. In addition, on $A_{2}$, observe that the number of centered points in the corrupted sample below $\xi_{1-\zeta/4,\tilde{X}}$ is at least $n(1-\zeta/2)$. As a result, (C.1) holds on $A_{2}$. An analogous argument yields (C.2) on $A_{2}$. Therefore, $E\subset A_{2}$.
Letting $A_{3}=A_{1}\cap A_{2}$, using the previous inequality, the definition of $\zeta$ we have that $A_{3}$ occurs with probability at least
$\begin{split}1-\frac{\delta}{8}-\frac{\delta}{8}=1-\frac{\delta}{4}.\end{split}$ (C.6)
We now show that, on $A_{3}$, the uncorrupted mean estimate is close to the true mean with high probability. Observe that on $A_{3}$,
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\tilde{\xi}_{\zeta},\tilde{\xi}_{1-\zeta}}(X_{i})$ $\leq\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})$
$\leq\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})+\frac{1}{n}\sum_{i=1}^{n}[\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})].$
Observe the following equality
$\mathbb{E}\phi_{\xi_{4\zeta,\tilde{X}},\xi_{1-\zeta/4,\tilde{X}}}(\tilde{X}_{1})$ $=\mathbb{E}(\xi_{4\zeta,\tilde{X}}\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\})+\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\})$
$\hskip 150.00023pt+\mathbb{E}(\tilde{X}\mathbbm{1}\left\{\xi_{4\zeta,\tilde{X}}<\tilde{X}<\xi_{1-\zeta/4,\tilde{X}}\right\})$
$=\mathbb{E}(\xi_{4\zeta,\tilde{X}}\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\})+\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\})$
$\hskip 150.00023pt-\mathbb{E}(\tilde{X}(\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\}+\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\}))$
$=\mathbb{E}((\xi_{4\zeta,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\})+\mathbb{E}((\xi_{1-\zeta/4,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\})$
$=\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})+\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})$
$=|\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})|-|\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})|.$
Now, we have that
$\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})$ $=\mu+\mathbb{E}\phi_{\xi_{4\zeta,\tilde{X}},\xi_{1-\zeta/4,\tilde{X}}}(\tilde{X}_{1})$
$=\mu+|\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})|-|\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})|$
$\leq\mu+\mathcal{E}(4\zeta,X),$
where $\mathcal{E}(a,X)=|\mathbb{E}(\xi_{a,\tilde{X}}\vee\tilde{X})|\vee|\mathbb{E}(\xi_{1-a,\tilde{X}}\wedge\tilde{X})|$. Similarly,
$\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})=\mu-|\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})|+|\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})|$ $\geq\mu-\mathcal{E}(\zeta/4,X).$
Now, let
$\overline{\omega}=\frac{1}{n}\sum_{i=1}^{n}\omega_{i}=\frac{1}{n}\sum_{i=1}^{n}[\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})]$
and observe that $\overline{\omega}$ is an independent sum of centered random variables, bounded above by $(\xi_{1-\zeta/4,\tilde{X}}+\mathcal{E}(\zeta/4,X))/n$. Moreover, $\text{var}(\omega_{i})\leq\sigma^{2}$. We then have that by Bernstein’s inequality
$\Pr\left(\overline{\omega}\geq t\right)$ $\leq\exp\left(-\frac{t^{2}/2}{\operatorname{Var}(\omega_{1})/n+(\xi_{1-\zeta/4,\tilde{X}}+\mathcal{E}(\zeta/4,X))t/3n}\right)$
$=\exp\left(-\frac{nt^{2}/2}{\operatorname{Var}(\omega_{1})+(\xi_{1-\zeta/4,\tilde{X}}+\mathcal{E}(\zeta/4,X))t/3}\right)$
$\leq\exp\left(-\frac{nt^{2}/2}{\sigma^{2}+(\xi_{1-\zeta/4,\tilde{X}}+\mathcal{E}(\zeta/4,X))t/3}\right).$
Thus, taking
$t$ $=\frac{\sqrt{[2\left(\mathcal{E}(\zeta/4,X)+\xi_{1-\zeta/4,\tilde{X}}\right)\log(1/p)]^{2}+72n\sigma^{2}\log(1/p)}+2\left(\mathcal{E}(\zeta/4,X)+\xi_{1-\zeta/4,\tilde{X}}\right)\log(1/p)}{6n}$
$=\frac{\sqrt{[\left(\mathcal{E}(\zeta/4,X)+\xi_{1-\zeta/4,\tilde{X}}\right)\log(1/p)]^{2}+18n\sigma^{2}\log(1/p)}+\left(\mathcal{E}(\zeta/4,X)+\xi_{1-\zeta/4,\tilde{X}}\right)\log(1/p)}{3n}$
renders $P(\overline{\omega}\geq t)\leq p$. Consider that, with
$t^{*}=\sigma\sqrt{\frac{6\log(1/p)}{n}}+\frac{2\left(\mathcal{E}(\zeta/4,X)+\xi_{1-\zeta/4,\tilde{X}}\right)\log(1/p)}{3n},$
we have $t\leq t^{*}$ and thus $P(\overline{\omega}\geq t^{*})\leq P(\overline{\omega}\geq t)\leq p$. Taking $p=\delta/4$, combined with (C.6) and the assumption that $\delta\geq 4e^{-n}$, yields that, with probability at least $21-\delta/4-\delta/4=1-\delta/$ 2, there exists universal constants $c,c^{\prime}>0$ such that
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{\zeta/4,X},\xi_{1-\zeta/4,X}}(X_{i})$ $\leq\mu+\mathcal{E}(4\zeta,X)+c\left(\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}+\frac{2\mathcal{E}(\zeta/4,X)\log(4/\delta)}{3n}\right)$
$\leq\mu+c^{\prime}\mathcal{E}(\zeta/4,X)+c\left(\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}\right).$
Applying Chebyshev’s inequality yields that
$\frac{\zeta}{4}=\Pr\left(\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right)\leq\frac{\sigma^{2}}{\xi_{1-\zeta/4,\tilde{X}}^{2}}\implies\xi_{1-\zeta/4,\tilde{X}}\leq\frac{\sqrt{4}\sigma}{\sqrt{\zeta}}.$
Since $\delta/4\geq e^{-n}$ then $\log(4/\delta)\leq n$, and so $\sqrt{\log(4/\delta)/n}\geq\log(4/\delta)/n$. Then, using the result from Chebyshev’s inequality, with probability at least $21-\delta/$ 2,
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mu$ $\lesssim\mathcal{E}(\zeta/4,X)+\sigma\sqrt{\frac{\log(4/\delta)}{n}}.$
Lastly, we bound $\mathcal{E}(\zeta/4,X)$ above. Using the triangle inequality, Holder’s inequality and Chebyshev’s inequality,
$|E((X-\xi_{1-a})\mathbbm{1}\left\{X>\xi_{1-a})\right\}|$ $\leq\mathbb{E}(|\tilde{X}\mathbbm{1}\left\{X>\xi_{1-a}\right\}|)+\mathbb{E}(|\xi_{1-a}|\mathbbm{1}\left\{X>\xi_{1-a}\right\})$
$\leq\sigma\Pr\left(X>\xi_{1-a}\right)^{1/2}+\sigma\sqrt{a}$
$\lesssim\sigma\sqrt{a}.$
Similarly, $|E((X-\xi_{a})\mathbbm{1}\left\{X<\xi_{a})\right\}|\lesssim\sigma\sqrt{a}$. Taken together, these two inequalities give $\mathcal{E}(\zeta/4,X)\lesssim\sigma\sqrt{\zeta}$. Thus, with probability at least $21-\delta/$ 2,
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mu$ $\lesssim\sigma\sqrt{\zeta}+\sigma\sqrt{\frac{\log(4/\delta)}{n}}.$
An analogous argument holds to give a similar lower bound on the same quantity.
Next, we consider
$\left|\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i}^{\prime})\right|,$
following the argument in Lugosi and Mendelson, [28]. Using Chebyshev’s inequality, note that the maximal gap for a single term in the summation,
$\left|\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i}^{\prime})\right|\leq|\xi_{1-\zeta/4,X}|+|\xi_{4\zeta,X}|\lesssim\sigma/\sqrt{\zeta}.$
Since $\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})=\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i}^{\prime})$ for all but $2\eta n$ (or fewer) points, we get that
$\left|\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i}^{\prime})\right|\lesssim\eta\sigma/\sqrt{\zeta}\lesssim\sigma\sqrt{\zeta}.$
The last inequality follows since, by assumption, $\eta\leq\zeta/16$.
Recall that
$\tilde{\mu}=\frac{1}{n}\sum_{i=1}^{n}\phi_{\tilde{\xi}_{\zeta},\tilde{\xi}_{1-\zeta}}(X_{i}^{\prime})+Z_{1}\frac{(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}}\leq\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{4\zeta},\xi_{1-\zeta/4}}(X_{i}^{\prime})+Z_{1}\frac{(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}}.$
Combining this with (C.6), the preceding tail argument, and the fact that $\Pr\left(Z_{1}\geq 1\right)=e^{-n}\leq\delta/4$, we have that with probability at least $1-\delta,$
$|\tilde{\mu}-\mu|\lesssim\sigma\sqrt{\zeta}+\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{Z_{1}(\tilde{\xi}_{1-\zeta}-\tilde{\xi}_{\zeta})}{n\varepsilon_{3}}\\ \lesssim\sigma\sqrt{\zeta}+\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{Z_{1}(\xi_{1-\zeta/4}-\xi_{\zeta/4})}{n\varepsilon_{3}}\\ \lesssim\sigma\sqrt{\zeta}+\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{\xi_{1-\zeta/4}-\xi_{\zeta/4}}{n\varepsilon_{3}}.$
It suffices to bound the term $\xi_{1-\zeta/4}-\xi_{\zeta/4}$. To this end, note that $\xi_{1-\zeta/4}-\xi_{\zeta/4}=\xi_{1-\zeta/4,X}-\xi_{\zeta/4,X}$. In addition, we have that $\zeta\gtrsim 1/n$ and so
$\xi_{1-\zeta/4}-\xi_{\zeta/4}\leq 2(\xi_{1-\zeta/4,X}\vee(-\xi_{\zeta/4}))\leq 2(\xi_{1-1/4n,X}\vee(-\xi_{1/4n,X})).$
Suppose that $\xi_{1-1/4n,X}>0$, then Chebyshev’s inequality implies that
$1/4n$ $=\Pr\left(X\geq\xi_{1-1/4n,X}\right)\leq\Pr\left(|X|\geq\xi_{1-1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1-1/4n,X}^{2}}\implies\xi_{1-1/4n,X}\lesssim\sigma\sqrt{n}.$
In addition, if $\xi_{1-1/4n,X}<0$, then $\xi_{1-1/4n,X}\lesssim\sigma\sqrt{n}$. Similarly, suppose that $\xi_{1/4n,X}<0$. Again, Chebyshev’s inequality implies that
$1/4n$ $=\Pr\left(X\leq\xi_{1/4n,X}\right)\leq\Pr\left(|X|\geq-\xi_{1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1/4n,X}^{2}}\implies-\xi_{1/4n,X}\lesssim\sigma\sqrt{n}.$
Thus, $\xi_{1-\zeta/4}-\xi_{\zeta/4}\lesssim\sigma\sqrt{n}.$
Appendix D Zero-concentrated DP quantiles (zCDP)
We now define a version of the estimator that satisfies a weaker notion of privacy called zero-concentrated DP. The algorithm proceeds as follows: Let $V,V_{1},V_{2},\ldots\sim\mathcal{N}(0,1)$. Given $\beta>1$, a lower bound $\ell$, data set $\{Y_{i}\}_{i=1}^{n}$ with empirical distribution $F_{n}$, and quantile $q>1/2$, we do the following:
  1. 1.
    First, compute $\hat{q}=q+V/n\sqrt{\rho}$.
  2. 2.
    While $F_{n}(\beta^{i}+\ell-1)+V_{i}/n\sqrt{\rho}<\hat{q}$, increase $i$ by one at a time. Otherwise, stop.
  3. 3.
    If we stop at $k$, then output $\beta^{k}+\ell-1$.
If instead $1/2>q\to 0$, we negate everything and take an upper bound as input instead of lower bound. The idea is that the upper bound is far from the minimum, etc. Call these estimates $\tilde{\xi}_{q,\mathbb{Y}_{n}}$. We will omit $\mathbb{Y}_{n}$ when it is obvious.
We first show that adding Gaussian noise with variance $1/n^{2}\rho$ to the queries in the private quantile estimation results in $\rho$-zCDP.
Lemma D.1.
For all $0\leq q\leq 1$ and $\rho_{1},\rho_{2}>0$, $\tilde{\xi}_{q}^{\prime}$ satisfies $(\rho_{1}+\rho_{2})$-zCDP.
Proof.
Consider neighboring datasets $\mathbf{x}_{n},\mathbf{x}_{n}^{\prime}$ with empirical distribution functions, $F_{n}$ and $F_{n}^{\prime}$, respectively. Let ${\rm D}_{\alpha}(P||Q)$ denotes the $\alpha$-Renyi Divergence between two measures $P$ and $Q$. That is,
${\rm D}_{\alpha}(P||Q)=\frac{1}{\alpha-1}\log\int\left(\frac{dP}{dQ}\right)^{\alpha}dQ.$
For two random variables on the same probability space, $X\sim P$ and $Y\sim Q$, let ${\rm D}_{\alpha}(X||Y)={\rm D}_{\alpha}(P||Q)$. The aim is to show that for any $\alpha>1$, $11/2<q<$ 1 and any pair of adjacent databases $\mathbf{x}_{n},\mathbf{x}_{n}^{\prime}$, we have that
${\rm D}_{\alpha}(\tilde{\xi}_{q,\mathbf{x}_{n}}||\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}})\leq\alpha\rho.$
To this end, let $K$ be the random integer at which the process stops on the original dataset $\mathbf{x}_{n}$, and let $K^{\prime}$ be the random integer at which the process stops on the neighboring dataset $\mathbf{x}_{n}^{\prime}$. It follows by post-processing, (see Lemma 15 of [8]) that ${\rm D}_{\alpha}(\tilde{\xi}_{q,\mathbf{x}_{n}}||\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}})\leq{\rm D}_{\alpha}(K||K^{\prime})$, since $\tilde{\xi}_{q,\mathbf{x}_{n}^{\prime}}=\beta^{K}+\ell-1$. It suffices to show that ${\rm D}_{\alpha}(K||K^{\prime})\leq\rho\alpha$ for all $\alpha>1$.
To this end, let $M\stackrel{{\scriptstyle\text{d}}}{{=}}K$, let $\nu_{j}$ be the joint distribution of the sequence of random variables $Z=(K,V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$ and let $\nu_{j}^{\prime}$ be the distribution of $Z^{\prime}=(K^{\prime},V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Take $I$ to be the function returning the first element, so that $I(Z)=K$ and $I(Z^{\prime})=K^{\prime}$. It follows by Lemma 15 of [8] (post-processing) that
${\rm D}_{\alpha}(K||K^{\prime})={\rm D}_{\alpha}(I(\nu_{j})||I(\nu_{j}^{\prime}))\leq{\rm D}_{\alpha}(\nu_{j}||\nu_{j}^{\prime}).$
We now consider ${\rm D}_{\alpha}(\nu_{j}||\nu_{j}^{\prime})$ for arbitrary $j$. Let $Q_{j}$ and $Q_{j}^{\prime}$ denote the distribution of $K$ and $K^{\prime}$ conditional on $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Let $\mu_{j}$ be the joint distribution of the sequence of random variables $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, we have that by Lemma 15 of [8] (composition),
${\rm D}_{\alpha}(\nu_{j}||\nu_{j}^{\prime})\leq\sup_{j\geq 0}[{\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})+{\rm D}_{\alpha}(\mu_{j}||\mu_{j})]=\sup_{j\geq 0}{\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime}).$
It then suffices to show that, for any $j\geq 0$, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho$. To this end, let $(v_{1}/n\sqrt{\rho},\ldots,v_{j-1}/n\sqrt{\rho},v_{j+1}/n\sqrt{\rho},\ldots)$ be the realizations of $(V_{1}/n\sqrt{\rho},\ldots,V_{j-1}/n\sqrt{\rho},V_{j+1}/n\sqrt{\rho},\ldots)$. Next, for any positive integer $k$, let $z_{k}=F_{n}(\beta^{k}+\ell-1)$, $\tau_{k}=\max_{i<k}[F_{n}(\beta^{i}+\ell-1)+v_{i}]$, $z_{k}^{\prime}=F^{\prime}_{n}(\beta^{k}+\ell-1)$ and $\tau_{k}^{\prime}=\max_{i<k}[F_{n}^{\prime}(\beta^{i}+\ell-1)+v_{i}]$. To prove the result, we first have that
$\Pr\left(K=k\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)\\ =\Pr\left(\tau_{k}\leq\hat{q}<z_{k}+V_{k}/n\sqrt{\rho}\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)\\ =\Pr\left(\tau_{k}\leq\hat{q}<z_{k}+V_{k}/n\sqrt{\rho}\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1}\right)\coloneqq b_{k}.$
Next, letting $\varphi(x)$ be the standard normal density, $q^{\prime}(x)=q+x/n\sqrt{\rho}+\tau_{k}^{\prime}-\tau_{k}$ and $\kappa(x)=x+n\sqrt{\rho}(\tau_{k}^{\prime}-\tau_{k}+z_{k}-z_{k}^{\prime})$ , by definition, it holds that
[…]
In the above integral, it is helpful to note that $x$ corresponds to the random variable $V$, and $y$ corresponds to the random variable $V_{k}$. Moreover, the indicator depends on $(V_{1},\dots,V_{k-1})$ only through $\tau_{k}$. Since $V$ and $V_{k}$ are both independent of $(V_{1},\dots,V_{k-1})$, the conditional densities are equal to the marginal densities. Continuing,
[…]
Next, we make the change of variable $u_{1}=\kappa(y)/n\sqrt{\rho}$ and $u_{2}\coloneqq u_{2}(x)=q^{\prime}(x)-q$.
$b_{k}$ $=(n^{2}\rho)^{-1}\int_{\mathbb{R}}\int_{\mathbb{R}}\mathbbm{1}\left\{\tau_{k}^{\prime}<u_{2}+q<u_{1}+z_{k}^{\prime}-\tau_{k}^{\prime}\right\}\varphi(n\sqrt{\rho}(u_{2}+\tau_{k}-\tau_{k}^{\prime}))$
×$\hskip 200.0003pt\times\varphi(n\sqrt{\rho}(u_{1}+\tau_{k}-z_{k}+z_{k}^{\prime}-\tau_{k}^{\prime}))du_{1}du_{2}$
$=\Pr\left(\tau_{k}^{\prime}<U_{2}+q<U_{1}+z_{k}^{\prime}\right)$
$=\Pr\left(\tau_{k}^{\prime}<U_{3}+q<z_{k}^{\prime}\right)\coloneqq h_{k}(U_{3}),$
where $U_{1}\sim\mathcal{N}(\tau_{k}^{\prime}-\tau_{k}-z_{k}^{\prime}+z_{k},(n^{2}\rho)^{-1}),U_{2}\sim\mathcal{N}(\tau_{k}^{\prime}-\tau_{k},(n^{2}\rho)^{-1})$ and $U_{3}=U_{2}-U_{1}$, with $U_{1}\perp U_{2}$. By definition, i.e., the properties of the normal distribution, we have that $U_{3}\sim\mathcal{N}(\tau_{k}^{\prime}-\tau_{k}-\tau_{k}^{\prime}+\tau_{k}+z_{k}^{\prime}-z_{k},2(n^{2}\rho)^{-1})=\mathcal{N}(z_{k}^{\prime}-z_{k},2(n^{2}\rho)^{-1})$. In addition, observe that by definition,
$\Pr\left(K^{\prime}=k\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$ $=\Pr\left(\tau_{k}^{\prime}\leq\hat{q}<z_{k}^{\prime}+V_{k}/n\sqrt{\rho}\Big{|}\{V_{i}=v_{i}\}_{i=1}^{k-1},\{V_{i}=v_{i}\}_{i=k+1}^{\infty}\right)$
$=h_{k}((n\sqrt{\rho})^{-1}(V-V_{j})).$
Using these facts yields that
${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})$ $=\frac{1}{\alpha-1}\sup_{j>0}\mathbb{E}_{M}\left[\left(\frac{\Pr\left(K=M\Big{|}\{V_{i}=v_{i}\}_{i=1}^{j-1},\{V_{i}=v_{i}\}_{i=j+1}^{\infty}\right)}{\Pr\left(K^{\prime}=M\Big{|}\{V_{i}=v_{i}\}_{i=1}^{j-1},\{V_{i}=v_{i}\}_{i=j+1}^{\infty}\right)}\right)^{\alpha-1}\right]$
$=\frac{1}{\alpha-1}\sup_{j>0}\mathbb{E}_{M}\left[\left(\frac{h_{j}(U_{3})}{h_{j}((n\sqrt{\rho})^{-1}(V-V_{j}))}\right)^{\alpha-1}\right]$
$\leq\frac{1}{\alpha-1}\sup_{j>0}\mathbb{E}_{M}\left[\left(\frac{U_{3}}{(n\sqrt{\rho})^{-1}(V-V_{j})}\right)^{\alpha-1}\right]$
$=\alpha 2\rho n^{2}\mathbb{E}(U_{3})^{2}/2$
$=\alpha\rho n^{2}\mathbb{E}(U_{3})^{2},$
where the second last line uses the fact that the Renyi divergence between two normal distributions with variance $\sigma^{2}$ and means $\mu_{1},\mu_{2}$ is given by $\frac{1}{2}\alpha(\mu_{1}-\mu_{2})^{2}/2\sigma^{2}.$ Now, we just need to bound the expectation $\mathbb{E}(U_{3})$, from which it follows by definition that $-1/n<\mathbb{E}(U_{3})<1/n$. As a result, we have that ${\rm D}_{\alpha}(Q_{j}||Q_{j}^{\prime})\leq\alpha\rho,$ which completes the proof. ∎
Appendix E Proving finite sample bounds for private quantiles
The following lemma gives a concentration result related to the purely differentially private quantile computed from any sample.
Lemma E.1.
Given the corrupted sample $Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$, and $t>0$, we have that for all $1/2<q\leq 1$, $\ell\leq\hat{\xi}_{q}$, and $1<\beta\leq\left.(\hat{\xi}_{q+t}-\ell+1)\right/(\hat{\xi}_{q-t}-\ell+1)$ it holds that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{\log\beta}+1\right)\exp\left(-nt(\varepsilon_{1}\wedge\varepsilon_{2})\right),$
and for all $20<q\leq 1/$ 2, $u\geq\hat{\xi}_{q}$, and $1<\beta\leq\left.(u-\hat{\xi}_{q-t}+1)\right/(u-\hat{\xi}_{q+t}+1)$ it holds that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{q+t}+1)}{\log\beta}+1\right)\exp\left(-nt(\varepsilon_{1}+\varepsilon_{2})\right).$
Proof of Lemma E.1.
We start by demonstrating the first inequality. Let $\hat{q}=q+V/n\varepsilon_{1}\geq q$. Observe that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)=\Pr\left(q-F_{n}(\tilde{\xi}_{q})>t\right)+\Pr\left(F_{n}(\tilde{\xi}_{q})-q>t\right)\coloneqq I+II.$
First, we must ensure that there is some $k\in\mathbb{N}$ such that $F_{n}(\beta^{k}-1+\ell)$ that is within $t$ of $q$. Let $k$ be the integer at which the process terminates. We have that
$q-F_{n}(\tilde{\xi}_{q})\geq t$ $\iff\beta^{k}-1+\ell\leq\hat{\xi}_{q-t}\iff k\leq\frac{\log(\hat{\xi}_{q-t}-\ell+1)}{\log\beta}\coloneqq k_{1}$
$F_{n}(\tilde{\xi}_{q})-q\geq t$ $\iff\beta^{k}-1+\ell\geq\hat{\xi}_{q+t}\iff k\geq\frac{\log(\hat{\xi}_{q+t}-\ell+1)}{\log\beta}\coloneqq k_{2}.$
Thus, as long as there is an integer between $k_{1}$ and $k_{2}$, there is a valid $k\in\mathbb{N}$. This is equivalent to saying $1\leq k_{2}-k_{1}$, giving
11 11 $\leq\frac{\log(\hat{\xi}_{q+t}-\ell+1)}{\log\beta}-\frac{\log(\hat{\xi}_{q-t}-\ell+1)}{\log\beta}\implies\beta\leq\frac{\hat{\xi}_{q+t}-\ell+1}{\hat{\xi}_{q-t}-\ell+1}.$
This holds by assumption, and so a valid $k$ must exist. We now consider $I$.
Using the definition of $k$, the fact that $V$ and $\{V_{i}\}_{i=1}^{\infty}$ are standard exponential random variables and a union bound, it holds that
$\Pr\left(k<k_{1}\right)$ $\leq\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\varepsilon_{2}\geq\hat{q}-F_{n}(\beta^{i}+\ell-1)\right)$
$\leq\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\varepsilon_{2}\geq\hat{q}-F_{n}(\beta^{k_{1}}+\ell-1)\right)$
$=\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\varepsilon_{2}\geq q+V/n\epsilon_{1}-F_{n}((\hat{\xi}_{q-t}-\ell+1)+\ell-1)\right)$
$\leq k_{1}\Pr\left(V_{k_{1}}-V\geq tn(\varepsilon_{1}\wedge\varepsilon_{2})\right)$
$=\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{2\log\beta}\exp\left(-nt(\varepsilon_{1}\wedge\varepsilon_{2})\right).$
Now, we consider $II$. Observe that
$II=\Pr\left(k>k_{2}\right)$ $=\prod_{i=1}^{k_{2}}\Pr\left(V_{i}/n\varepsilon_{2}+F_{n}(\beta^{i}+\ell-1)\leq\hat{q}\right)$
$\leq\Pr\left(V_{i}/n\varepsilon_{2}+F_{n}(\beta^{k_{2}}+\ell-1)\leq\hat{q}\right)$
$=\Pr\left(V_{i}/n\varepsilon_{2}+F_{n}((\hat{\xi}_{q+t}-\ell+1)+\ell-1)\leq q+V/n\epsilon_{1}\right)$
$\leq\Pr\left(tn(\varepsilon_{2}\wedge\varepsilon_{2})\leq V-V_{i}\right)$
$\leq\frac{1}{2}\exp\left(-nt(\varepsilon_{1}\wedge\varepsilon_{2})\right).$
Then, for the second inequality, since $0<q\leq\frac{1}{2}$, the algorithm is run with $\ell=-u$, quantile $1-q$, and each data point negated, defining $Y=-X$. Then $\tilde{\xi}_{q,X}=-\tilde{\xi}_{q^{\prime},Y}$, $\hat{\xi}_{q,X}=-\hat{\xi}_{q^{\prime},Y}$, and $F_{n,X}(k)=1-F_{n,Y}(-k)$.
Note that $0<q\leq\frac{1}{2}\iff\frac{1}{2}<q^{\prime}\leq 1$, $u\geq\hat{\xi}_{q,X}\iff\ell\leq\hat{\xi}_{q^{\prime},Y}$, and
$1\leq\beta\leq\frac{u-\hat{\xi}_{q-t,X}+1}{u-\hat{\xi}_{q+t,X}+1}\iff 1\leq\beta\leq\frac{-\ell+\hat{\xi}_{1-(q-t),Y}+1}{-\ell+\hat{\xi}_{1-(q+t),Y}+1}\iff 1\leq\beta\leq\frac{\hat{\xi}_{q^{\prime}+t,Y}-\ell+1}{\hat{\xi}_{q^{\prime}-t,Y}-\ell+1}.$
Thus, the first inequality can be applied with $\ell=-u$, $q^{\prime}$, and $Y$. This gives
$P(|q^{\prime}-F_{n,Y}(\tilde{\xi}_{q^{\prime}-t,Y})|>t)\leq\left(\frac{\log(\hat{\xi}_{q^{\prime},Y}+1-\ell)}{\log(\beta)}+1\right)\exp\left(-nt(\epsilon_{1}\wedge\epsilon_{2})\right).$
Note that
$|q^{\prime}-F_{n,Y}(\tilde{\xi}_{q^{\prime},Y})|=|1-q-(1-F_{n,X}(-\tilde{\xi}_{q^{\prime},Y}))|=|F_{n,X}(\tilde{\xi}_{q,X})-q|.$
For the upper bound, consider that
$\frac{\log(\hat{\xi}_{q^{\prime},Y}+1-\ell)}{\log(\beta)}=\frac{\log(-\hat{\xi}_{1-(1-q-t),X}+1+u)}{\log(\beta)}=\frac{\log(u-\hat{\xi}_{q+t,X}+1)}{\log(\beta)}.$
This completes the proof. ∎
The following lemma gives a concentration result related to the zero-concentrated differentially private quantile computed from any sample.
Lemma E.2.
Given the corrupted sample $Y_{1}^{\prime},\ldots,Y_{n}^{\prime}$, and $t>0$, we have that for all $1/2<q\leq 1$, $\ell\leq\hat{\xi}_{q}$, and $1<\beta\leq(\hat{\xi}_{q+t}-\ell+1)/(\hat{\xi}_{q-t}-\ell+1)$, we have that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{\log\beta}+1\right)\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right),$
and for all $20<q\leq 1/$ 2, $u\geq\hat{\xi}_{q}$, and $1<\beta\leq(u-\hat{\xi}_{q-t}+1)/(u-\hat{\xi}_{q+t}+1)$ it holds that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{q+t}+1)}{\log\beta}+1\right)\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right).$
Proof of Lemma E.2.
The proof mirrors that of Lemma E.1. We start by demonstrating the first inequality. Let $\hat{q}^{\prime}=q+V/n\sqrt{\rho}_{1}$. Observe that
$\Pr\left(|q-F_{n}(\tilde{\xi}_{q})|>t\right)=\Pr\left(q-F_{n}(\tilde{\xi}_{q})>t\right)+\Pr\left(F_{n}(\tilde{\xi}_{q})-q>t\right)\coloneqq I+II.$
First, we must ensure that there is some $k\in\mathbb{N}$ such that $F_{n}(\beta^{k}-1+\ell)$ that is within $t$ of $q$. Let $k$ be the integer at which the process terminates. We have that
$q-F_{n}(\tilde{\xi}_{q})\geq t$ $\iff\beta^{k}-1+\ell\leq\hat{\xi}_{q-t}\iff k\leq\frac{\log(\hat{\xi}_{q-t}-\ell+1)}{\log\beta}\coloneqq k_{1}$
$F_{n}(\tilde{\xi}_{q})-q\geq t$ $\iff\beta^{k}-1+\ell\geq\hat{\xi}_{q+t}\iff k\geq\frac{\log(\hat{\xi}_{q+t}-\ell+1)}{\log\beta}\coloneqq k_{2}.$
Thus, as long as there is an integer between $k_{1}$ and $k_{2}$, there is a valid $k\in\mathbb{N}$. This is equivalent to saying $1\leq k_{2}-k_{1}$, giving
11 11 $\leq\frac{\log(\hat{\xi}_{q+t}-\ell+1)}{\log\beta}-\frac{\log(\hat{\xi}_{q-t}-\ell+1)}{\log\beta}\implies\beta\leq\frac{\hat{\xi}_{q+t}-\ell+1}{\hat{\xi}_{q-t}-\ell+1}.$
This holds by assumption, and so a valid $k$ must exist. We now consider $I$.
Using the definition of $k$, the fact that $V$ and $\{V_{i}\}_{i=1}^{\infty}$ are standard normal random variables and a union bound, it holds that
$\Pr\left(k<k_{1}\right)$ $\leq\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\sqrt{\rho_{2}}\geq\hat{q}-F_{n}(\beta^{i}+\ell-1)\right)$
$\leq\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\sqrt{\rho_{2}}\geq\hat{q}^{\prime}-F_{n}(\beta^{k_{1}}+\ell-1)\right)$
$=\sum_{i=1}^{k_{1}}\Pr\left(V_{i}/n\sqrt{\rho_{2}}\geq q+V/n\sqrt{\rho_{1}}-F_{n}((\hat{\xi}_{q-t}-\ell+1)+\ell-1)\right)$
$\leq k_{1}\Pr\left(V_{k_{1}}-V\geq tn\sqrt{(\rho_{1}\wedge\rho_{2})}\right)$
$\leq\frac{\log(\hat{\xi}_{q-t}+1-\ell)}{2\log\beta}\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right).$
Here we use an exponential tail bound for the standard Gaussian, as we will be applying this result when $t<1$. Now, we consider $II$. Observe that
$II=\Pr\left(k>k_{2}\right)$ $=\prod_{i=1}^{k_{2}}\Pr\left(V_{i}/n\sqrt{\rho_{2}}+F_{n}(\beta^{i}+\ell-1)\leq\hat{q}\right)$
$\leq\Pr\left(V_{i}/n\sqrt{\rho_{2}}+F_{n}(\beta^{k_{2}}+\ell-1)\leq\hat{q}\right)$
$=\Pr\left(V_{i}/n\sqrt{\rho_{2}}+F_{n}((\hat{\xi}_{q+t}-\ell+1)+\ell-1)\leq q+V/n\sqrt{\rho_{1}}\right)$
$\leq\Pr\left(tn\sqrt{(\rho_{2}\wedge\rho_{2})}\leq V-V_{i}\right)$
$\leq\exp\left(-n(t\vee t^{2})(\sqrt{\rho_{1}\vee\rho_{1}^{2}}\wedge\sqrt{\rho_{2}\vee\rho_{2}^{2}})/4\right).$
Then, for the second inequality, the same argument as in Lemma E.1 applies. ∎
Appendix F Error in the proof of Theorem 1 of Lugosi and Mendelson, [28]
Here, we demonstrate that there is an error in the proof of Theorem 1 of Lugosi and Mendelson, [28]. Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\})\text{ and }\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, [28]. We show they do not hold by counterexample. For the remainder of this section, we use the notation of Lugosi and Mendelson, [28].
F.1 The lower bound
Using the notation in the paper, on page 401 at the very top, it states
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(X)\geq\mu-\mathbb{E}[\overline{X}\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})}].$
This inequality does not hold. Observe that, using the fact that $\mathbb{E}(\overline{X})=0$,
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(X)-\mu$ $=Q_{2\epsilon}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})+Q_{1-\epsilon/2}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})$
$\hskip 100.00015pt+\mathbb{E}(\overline{X}(1-\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}-\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})}))$
$=Q_{2\epsilon}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})+Q_{1-\epsilon/2}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})$
$\hskip 100.00015pt-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})$
$=-\mathbb{E}(\overline{X}\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})+Q_{2\epsilon}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})$
$\hskip 100.00015pt+\mathbb{E}((Q_{1-\epsilon/2}(\overline{X})-\overline{X})\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})}).$
It then suffices to find some $F$ such that
$Q_{2\epsilon}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})+\mathbb{E}((Q_{1-\epsilon/2}(\overline{X})-\overline{X})\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})<0.$
The term $\mathbb{E}((Q_{1-\epsilon/2}(\overline{X})-\overline{X})\mathbbm{1}_{\overline{X}\geq Q_{1-\epsilon/2}(\overline{X})})\leq 0$ for all $F$. Then, for example, standard normal distribution has $Q_{2\epsilon}(\overline{X})\mathbb{E}(\mathbbm{1}_{\overline{X}\leq Q_{2\epsilon}(\overline{X})})<0$, so the bound cannot hold in general.
F.2 The upper bound
We now give a counterexample to
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(X)\leq\mu+\mathbb{E}[\overline{X}\mathbbm{1}_{\overline{X}>Q_{1-\epsilon/2}(\overline{X})}],$ (F.1)
for all $\epsilon<1/2$, $E(\overline{X})=0$ and $F$ such that $Var(\overline{X})<\infty$. Consider a random variable $\overline{X}\sim F$, where $F$ is a discrete distribution with four points, $x_{1}<x_{2}<0<x_{3}<x_{4}$. Suppose that $\Pr\left(X=x_{i}\right)=p_{i}$ for $i\in[4]$ and suppose that $\overline{X}$ has mean 0: $\mathbb{E}(\overline{X})=\sum_{i=1}^{4}x_{i}p_{i}=0$. If $x_{i}$ are finite, then the variance of $\overline{X}$ is finite. Therefore, $F$ is in the class of considered distributions. (Continuous analogues of this example can be derived, by placing concentrated density in a neighbourhood of $x_{i}$ for each $x_{i}$.) First, for $p_{1}<2\epsilon<p_{2}+p_{1}$ and $p_{1}+p_{2}<1-\epsilon/2\iff\epsilon<2(p_{4}+p_{3})$, it holds that
$\mathbb{E}\phi_{\mu+Q_{2\epsilon}(\overline{X}),\mu+Q_{1-\epsilon/2}(\overline{X})}(\overline{X})$ $=(p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3}$
$\mathbb{E}[\overline{X}\mathbbm{1}_{\overline{X}>Q_{1-\epsilon/2}(\overline{X})}]$ $=x_{4}p_{4}.$
To prove that (F.1) does not hold for $F$, we must show that $((p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3})>x_{4}p_{4}$ for appropriate $p_{i}$ and $x_{i}$. So we just need a solution to the set of constraints:
$2\displaystyle 0<\epsilon<1/$ 2
$\epsilon<2(p_{4}+p_{3})$
$p_{1}<2\epsilon<p_{2}+p_{1}$
$x_{1}<x_{2}<0<x_{3}<x_{4}$
$p^{\prime}1_{4}=1$
$x^{\prime}p=0$
$1\displaystyle 0<p_{i}<$ 1
$((p_{1}+p_{2})x_{2}+(p_{3}+p_{4})x_{3})>x_{4}p_{4}.$
Taking $p_{1}=0.01$,$p_{2}=0.6$,$p_{3}=0.38$,$p_{4}=0.01$, $\epsilon=0.05$, $x_{1}=-57.1$,$x_{2}=-1$, $x_{3}=3$,$x_{4}=3.1$ gives a solution. You can get a similar counterexample taking $F$ to be a continuous distribution which is a mixture of uniform densities on the neighborhoods $x_{i}\pm\delta$ for some small $\delta>0$, with weights $p_{1},\ldots,p_{4}$. (You must also adjust $\epsilon$ by $\delta$).
Appendix G Skewed distributions and trimmed means
Here, we show that the non-private trimmed mean is inconsistent when $m/n\to p>0$ as $n\to\infty$ and the distribution is exponential. It follows by Slutsky’s theorem that the private trimmed mean is also inconsistent. (The proof can easily be extended to any skewed distribution, but it is simpler to present a concrete case.)
Lemma G.1.
Suppose that $\eta=0$ and $F$ is the standard exponential distribution. Then, when $m/n\to p>0$ as $n\to\infty$, the non-private trimmed mean is inconsistent.
We have that, by the law of large numbers, the (non-private) trimmed mean converges in probability to $\mathbb{E}(X|X\in[-\log(1-p),-\log(p)])$ as $n\to\infty$, where $X\sim Exp(1)$. Observe that for $0<a<b<\infty$, it holds that
[…]
$=-be^{-b}+ae^{-a}-e^{-b}+e^{-a}$
$=(a+1)e^{-a}-(b+1)e^{-b}.$
It follows that
$\mathbb{E}(X|X\in[-\log(1-p),-\log(p)])=(1-\log(1-p))(1-p)-(1\log p)p\neq 1.$
Therefore, for any non-vanishing clipping parameter $m=\lfloor pn\rfloor$, the trimmed mean is inconsistent. It follows by Slutsky’s theorem that the private trimmed mean is also inconsistent.